### 02 Apr 2026 v2

# Substitution for minimizing/maximizing a tropical linear (fractional) programming

Abstract
Tropical polyhedra seem to play a central role in static analysis of softwares. These tropical geometrical objects play also a central role in parity games especially mean payoff games and energy games. And determining if an initial state of such game leads to win the game is known to be equivalent to solve a tropical linear optimization problem. This paper mainly focus on the tropical linear minimization problem using a special substitution method on the tropical cone obtained by homogenization of the initial tropical polyhedron. But due to a particular case which can occur in the minimization process based on substitution we have to switch on a maximization problem. Nevertheless, forward-backward substitution is known to be strongly polynomial. The special substitution developed in this paper inherits the strong polynomiality of the classical substitution for linear systems. This special substitution must not be confused with the exponential execution time of the tropical Fourier-Motzkin elimination. Tropical fractional minimization problem with linear objective functions is also solved by tropicalizing the Charnes-Cooper’s transformation of a fractional linear program into a linear program developed in the usual linear algebra. Let us also remark that no particular assumption is made on the polyhedron of interest.
Finally, the substitution method is illustrated on some examples borrowed from the litterature.
Keywords: $(\max,+)$-algebra, tropical linear programming, fractional programming, algorithmic complexity.
MSC: 15A80, 16Y60, 90C32, 03D15.
1 Main notations and definitions
For every $\leq k^{\prime}, k,k$$k^{\prime}\in\mathbb{Z}$ we define the discrete interval $|k,k^{\prime}|]:=\{k,k+1,\ldots,k^{\prime}-1,k^{\prime}\}$. The inclusion of sets is denoted $\subseteq$ and the strict inclusion of set is denoted $\subset$.
Let us introduce the following notations and definitions for the $(\max,+)$-algebra. The reader is invinted to read eg. [3] to find more details on $(\max,+)$-algebra and idempotent semi-rings.
The $(\max,+)$-algebra denotes the set $\mathbb{R}_{\mbox{{\scriptsize$\mathbb{O}$}}}:=\mathbb{R}\cup\{-\infty\}$ equipped the “addition”: $(a,b)\mapsto\max(a,b)$ and the “multiplication” $(a,b)\mapsto a+b$.
In the sequel we will use the following notations: $a\oplus b:=\max(a,b)$, $a\otimes b:=a+b$ (sometimes $a\otimes b$ could be denoted $ab$). $\mathbbm{1}:=0$ is the neutral element for $\otimes$, $\mbox{{\scriptsize$\mathbb{O}$}}:=$-$\infty$ is the neutral element for $\oplus$, $\mathbb{O}$ is also the absorbing element of $\otimes$. We will use the following power notation:
$a^{\otimes(b)}:=a\times b\mbox{ ($\times$: usual multiplication)}.$
Note that in particular we have: $a^{\otimes(0)}=\mathbbm{1}$.
The product $a\otimes b^{\otimes(-1)}$ will also be denoted: $a/b$. Because division is required in this paper we have: $\mbox{{\scriptsize$\mathbb{O}$}}^{\otimes(-1)}=+\infty$. Thus, we work in the complete dioid by adding the top element $\infty$:
$(\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}},\oplus,\otimes)$ (1a)
with:
$\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}}:=\mathbb{R}\cup\{-\infty,+\infty\},$ (1b)
and the operations $\oplus$ and $\otimes$ are extended as follows:
$\forall a\in\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}},\forall b\neq\mbox{{\scriptsize$\mathbb{O}$}},\;a\oplus+\infty=+\infty,\;b\otimes+\infty=+\infty,\;\mbox{{\scriptsize$\mathbb{O}$}}\otimes+\infty=\mbox{{\scriptsize$\mathbb{O}$}}.$ (1c)
$\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}}$ is equipped with the natural order $\leq$ defined by:
$a\leq b\Leftrightarrow a\oplus b=b.$
For any matrix $A$, $a_{i,.}$ denotes the $i$th row of $A$, $a_{.,j}$ denotes the $j$th column of $A$ and $a_{i,j}$ denotes the entry $(i,j)$ of $A$.
The scalar operations are generalized to matrices as follows:
  • addition of two matrices: $(a_{i,j})\oplus(b_{i,j}):=(a_{i,j}\oplus b_{i,j})$
  • product of two matrices: the entry $(i,j)$ of the matrix $C=A\otimes B$ is defined by:
    $c_{i,j}:=\oplus_{k}a_{i,k}\otimes b_{k,j}$ ($\max_{k}(a_{i,k}+b_{k,j})$).
  • comparaison of two matrices: $A\leq B$ means $\forall i,j:a_{i,j}\leq b_{i,j}$.
  • all vectors are column vectors. And the transpose operator is denoted $(\cdot)^{\intercal}$.
  • for every $m\times n$-matrix $A$ we define the submatrix of $A$ denoted $A_{IJ}$ by:
    $A_{IJ}:=(a_{i,j})_{i\in I,j\in J}.$ (2)
    Where $I\subseteq[|1,m|]$ and $J\subseteq[|1,n|]$.
We will use the bold symbols for the following particular vectors and matrices.
[…]
  • The matrix $\boldsymbol{I}_{n}$ denotes the $n\times n$-identity matrix. Its diagonal entries are all $\mathbbm{1}$ and its off-diagonal entries are all $\mathbb{O}$.
  • $\boldsymbol{O}_{m,n}$ will denote the $m\times n$-null matrix (ie, all entries are $\mathbb{O}$). When $m=n$ the the square null matrix $\boldsymbol{O}_{n,n}$ is denoted $\boldsymbol{O}_{n}$.
To any $n\times n$-matrix $A$ we can associate a valued graph $G(A)$ with valuation $v$ such that the set of the vertices of $G(A)$ is $\{1,\ldots,n\}$. The valuation is defined by $v_{i,j}=a_{j,i}$ for all vertices $i,j$. We say that an arc $i\rightarrow j:=(i,j)$ exists if $v_{i,j}\neq\mbox{{\scriptsize$\mathbb{O}$}}$. A path $p$ of lenght $k\geq 1$ is a series of $k$ consecutive arcs which has the form: $p=i_{0}\rightarrow i_{1}\rightarrow\cdots i_{k-1}\rightarrow i_{k}$. And the weight of $p$ is defined by: $w(p):=v_{i_{0},i_{1}}\otimes\cdots\otimes v_{i_{k-1},i_{k}}=a_{i_{1},i_{0}}\otimes\cdots\otimes a_{i_{k},i_{k-1}}$. An elementary circuit of lenght $k\geq 1$ is a special path such that $i_{k}=i_{0}$ and $\forall l=1,\ldots,k-1$: $i_{l}$ appears only once.
Let $q,s$ be two $n$-dimensional vectors and $x=(x_{j})_{j=1}^{n}$. Then, to any inequality of the form:
$q^{\intercal}\otimes x\geq s^{\intercal}\otimes x$ (3a)
Noticing that (3a) is expressed in usual notations as $(\max(q_{1}+x_{1},\ldots,q_{n}+x_{n})\geq(\max(s_{1}+x_{1},\ldots,s_{n}+x_{n})$ one can associate the following equivalent disjunction recalling that $\max(a,b)\geq c\Leftrightarrow(a\geq c)\mbox{ or }(b\geq c)$:
$\textsf{OR}_{j=1}^{n}\{q_{j}\otimes x_{j}\geq s^{\intercal}\otimes x\}.$ (3b)
And one can associate the following equivalent conjonction recalling that $\max(a,b)\leq c\Leftrightarrow(a\leq c)\mbox{ and }(b\leq c)$:
$\textsf{AND}_{j=1}^{n}\{s_{j}\otimes x_{j}\leq q^{\intercal}\otimes x\}.$ (3c)
2 Introduction
2.1 Problem description
Let $A$ and $C$ be two $m\times n$-matrices and $b$ and $d$ be two $m$-dimensional column vectors. In this paper we aim to solve by a special substitution method the following max-plus linear programming problem called minimum problem research (mpr):
$\textsf{min}\{z=c^{\intercal}\otimes x,x\in\mathcal{P}(A,b,C,d)\cup\{\boldsymbol{\pm\infty}\}\},$ (4a)
where $\mathcal{P}(A,b,C,d)$ denotes the max-plus polyhedron defined by:
$\mathcal{P}(A,b,C,d):=\{x\in\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}}^{n}:A\otimes x\oplus b\geq C\otimes x\oplus d\},$ (4b)
But instead of considering the polyhedron we will develop substitution method on the max-plus cone $\mathcal{C}(A,b,C,d)$ obtained by introducing $h$ as the homogenization variable, ie.:
$\mathcal{C}(A,b,C,d):=\{(x,h)\in\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}}^{n+1}:A\otimes x\oplus b\otimes h\geq C\otimes x\oplus d\otimes h\}\cup\{\boldsymbol{\pm\infty}\}.$ (5a)
The linear system of inequalities $A\otimes x\oplus b\otimes h\geq C\otimes x\oplus d\otimes h$ is supposed to verify the following condition:
$\begin{array}[]{l}\forall i\in[|1,m|]\\ \{i^{\prime}:i^{\prime}\neq i\mbox{ and }\exists\alpha_{i^{\prime}}\in\mathbb{R}\mbox{ s.t. }a_{i,.}\otimes x\oplus b_{i}\otimes h\geq\alpha_{i^{\prime}}\otimes(c_{i^{\prime},.}\otimes x\oplus d_{i^{\prime}}\otimes h)\}=\emptyset.\end{array}$ (5b)
Remark 2.1
The condition (5b) can be assumed to be always true because $f\geq g_{1}$ and $f\geq g_{2}$ is equivalent to $f\geq g_{1}\oplus g_{2}$.
To formulate the minimization problem mpr first we borrow the Fourier’s trick (see eg. [13], [20]) used in the usual linear algebra and applying it to the $(\max,+)$-algebra. It comes that $\min(z=c^{\intercal}\otimes x\oplus c_{h}\otimes h)$ can be replaced by: $\min(z)\mbox{ and }z\geq c^{\intercal}\otimes x\oplus c_{h}\otimes h$. And then, we will solve the following minimization problem mpr defined as:
$\begin{array}[]{l}\min(z)\end{array}$ (6a)
such that:
$(x,h)\in\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}}^{n+1},$ (6b)
and
$\begin{array}[]{l}z\geq c^{\intercal}\otimes x\oplus c_{h}\otimes h\\ A\otimes x\oplus b\otimes h\geq C\otimes x\oplus d\otimes h.\end{array}$ (6c)
Where the homogenization variable $h$ satisfies the following condition:
$h>\mbox{{\scriptsize$\mathbb{O}$}}.$ (6d)
And
$\mbox{$z$ and $h$ are not substituable}.$ (6e)
Finally, among the variables of the problem $z,x_{1},\ldots,x_{n}$ we have:
[…]
Sometimes it will be useful to denote by $\textsf{mpr}(A,b,C,d,c,c_{h},z,x,h,(m,n))$ the mpr defined (6a)-(6b)-(6c)-(6d)-(6e)-(6f).
2.2 Motivations
Motivations for studying tropical linear programms have been widely explained in several previous works by several authors. In [5] author motivate tropical linear optimization by aiming to solve the problem of the synchronization of multiprocessors interactive systems. In [10] and [11] authors indicate further motivations to solve not only tropical linear programms but also fractional tropical linear programms (see § 5 for the description of such problem). One can mention eg.: static analysis of softwares, finding winning strategy in Mean Payoff Game.
2.3 Organization of the paper
Section 3 deals with the necessary materials to solve the minimization problem mpr. The substitution method is described in Section 4.
The reader must be aware that the substitution must not be confused with the tropical Fourier-Motzkin elimination procedure which has an exponential execution time [2].
The substitution is based on the prametrized research domain (22a)-(22b) for the $x_{j}$’s and the homogenization variable $h$. Dominating variable is defined in Definition 4.1. And the hierarchy $\preceq_{var}$ between variables $x_{j}$ (33) characterizes the fact that the value domain of the dominating variables has non trivial upper and lower bounds. The partition of the $(x,h)$-linear functions into $h$-bounded and $h$-unbounded functions is defined in Definition 4.2. The hierarchy $\preceq_{fct}$ between $(x,h)$-linear functions (36) is based on the obvious property that value domain of a $h$-bounded $(x,h)$-linear function $\alpha^{\intercal}\otimes x\oplus\beta\otimes h$ only depends on $\beta$ and is strictly included in the value domain of the function $\alpha^{\intercal}\otimes x$ (see (34a)-(34b)-(34c)). The Theorem 4.3 of § 4.2.2 characterizes the relevant variable $x_{j^{*}}$ which can be substituted. And the Theorem 4.2 of § 4.2.2 ensures the optimality of the result. The update of the mpr is described in § 4.2.3. Section 5 deals with the fractional linear programming. The method is illustrated on numerical examples in Section 6. Finally, we conclude in Section 7 where the complexity of our substitution method (about $\mathcal{O}(n^{2}m)$) is more discussed. We add some remarks on this work vs other works. Finally, the maximization problem is presented in appendix A.
2.4 The switch to a maximizing problem can occur
The appendix A dealing with maximization problem is needed because the following situation can occur in the substitution process for mpr. The reader has noticed that Fourier’s trick induces that the cost function is a lower bound of the variable $z$. Thus, because all linear functions are non-decreasing the research of lower bounds on the $x_{j}$’s is the priority. But in the substitution process we can have the following situation:
  • the cost function only depends on the homogenization variable $h$, ie. $z\geq c_{h}\otimes h$,
  • not all the variables have been substituted and they have no lower bounds but upper bounds exist at least for one variable,
  • the optimality is not ensured, ie. Theorem 4.2 of § 4.2.2 does not apply.
Noticing that minimazing $c_{h}\otimes h$ or maximazing $c_{h}\otimes h$ over the set of the remaining variables does not make differences one can switch to the maximization problem with $z\leq c_{h}\otimes h$ (Fourier’s trick for $\max$) and compute upper bounds on the remaining $x_{j}$’s of the problem.
Dually, the switch from maximization problem to minimization problem can occur.
3 Preliminary results
Let us consider two $m\times n$-matrices $L$ and $W$. And let us consider the cone $\mathcal{C}(L,W):=\{x:L\otimes x\geq W\otimes x\}$. To this cone we associate the function $\textsf{setrowtozero}(L,W)$ defined by:
[…]
3.1 Elementary results from interval arithmetic in dioids
In this subsection we present elementary materials dealing with interval arithmetic in the $(\max,+)$-algebra. The interval arithmetic on dioids has been already used in the context of automatic-control problem (see eg. [15], [14]). An interval is a set denoted $I:=[u,v]$ where $u,v\in\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}}$ and $u\leq v$. We define hereafter the elementary needed in this paper:
  • Addition of two intervals. Let $I_{i}=[u_{i},v_{i}]$, $i=1,2$ be two intervals, then the addition of $I_{1}$ and $I_{2}$ is denoted $I_{1}\oplus I_{2}$ and defined by:
    $u_{1},v_{1}]\oplus[u_{2},v_{2}]:=[u_{1}\oplus u_{2},v_{1}\oplus v_{2}].$ (8)
  • Scalar multiplication. Let $I=[u,v]$ be an interval and let $\alpha\in\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}}$. Then, the scalar multiplication of $I$ by the scalar $\alpha$ is denoted $\alpha\otimes I$ and defined by:
    $\alpha\otimes[u,v]:=[\alpha\otimes u,\alpha\otimes v].$ (9)
  • And the linear combination of $k$ intervals is the following interval defined as:
    $\oplus_{i=1}^{k}\alpha_{i}\otimes[u_{i},v_{i}]:=[\oplus_{i=1}^{k}\alpha_{i}\otimes u_{i},\oplus_{i=1}^{k}\alpha_{i}\otimes v_{i}].$ (10)
    Where $\alpha_{i}\in\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}}$, $i=1,\ldots,k$.
We will have to use results dealing with particular intervals. These particular intervals are defined as follows:
$\mathcal{I}:=\{[u,+\infty],u\in\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}}\}.$ (11)
Let us point out that the familly of intervals $\mathcal{I}$ is obviously stable by $(\max,+)$-linear combinations.
Proposition 3.1
Let $u_{l},+\infty]$, $l=1,\ldots,k$, $k\geq 2$ be $k$ intervals of $\mathcal{I}$ with $u_{1}\leq\ldots\leq u_{k}$. Then, we have the following series of intervals inclusion:
$u_{k},+\infty]\subseteq[u_{k-1},+\infty]\subseteq\cdots\subseteq[u_{1},+\infty].$ (12)
If Min denotes the minimum in the sense of interval inclusion then, from Proposition 3.1 we have the following noticeable interval equality:
$\cap_{l=1}^{k}[u_{l},+\infty]=\textsf{Min}\{[u_{l},+\infty],l=1,\ldots,k\}.$ (13)
Let $A$ be a $n\times n$-matrix, $b$ a $n$-dimensional vector and $x=(x_{i})_{i=1}^{n}$. We have the following well-known result.
Result 3.1 (Sub and super fix point)
Let us consider the following sets defined by:
$\mathcal{S}^{\leq}(A,b):=\{x:x\leq A\otimes x\oplus b\}.$ (14a)
and
$\mathcal{S}^{\geq}(A,b):=\{x:x\geq A\otimes x\oplus b\}.$ (14b)
If the following condition holds:
$\lim_{k\rightarrow+\infty}A^{\otimes(k)}=\boldsymbol{O}_{n}\mbox{ ($\boldsymbol{O}_{n}$: $n\times n$-null matrix)}$ (15)
then the greatest element of the set $\mathcal{S}^{\leq}(A,b)$ (resp. the smallest element of the set $\mathcal{S}^{\geq}(A,b)$) is:
$x=A^{*}\otimes b,$
where $A^{*}:=\boldsymbol{I}_{n}\oplus A\oplus A^{\otimes(2)}\oplus A^{\otimes(3)}\oplus\cdots$ is the infinite “sum” of the powers of the matrix $A$ known as the Kleene star of the matrix $A$. And under condition (15) we have: $A^{*}=\boldsymbol{I}_{n}\oplus A\oplus A^{\otimes(2)}\oplus\cdots\oplus A^{\otimes(n-1)}$.
This greatest element is the solution of the $(\max,+)$-linear equation: $x=A\otimes x\oplus b$ which is obtained by the saturation of the inequality: $x\leq A\otimes x\oplus b$.
There exist different conditions such that (15) is true. Notably the one which states that all elementary circuits of the valued graph $G(A)$ associated with the matrix $A$ have weight $\mathbbm{1}$.
We specify the previous Result 3.1 in the following case.
Proposition 3.2 (Valid inequality)
Let $a$ be a scalar, $a\in\mathbb{R}_{\mbox{{\scriptsize$\mathbb{O}$}}}$. Let $v$ be a $n$-dimensional vector. And let $y=(y_{i})_{i=1}^{n}$ be a $n$-dimensional column vector of variables in $\mathbb{R}_{\mbox{{\scriptsize$\mathbb{O}$}}}$. Then, an inequality of the form:
$a\otimes y_{j}\lesseqgtr v^{\intercal}\otimes y,$ (16)
is said to be valid iff the following condition is fulfilled:
$\mbox{$a\neq\mbox{{\scriptsize$\mathbb{O}$}}$ and $a>v_{j}$}.$ (17)
And in this case the inequality (16) is equivalent to:
$y_{j}\lesseqgtr a^{\otimes(-1)}\otimes\overline{v}^{\intercal}\otimes y,$ (18a)
where $\overline{v}$ is the $n$-dimensional vector defined by:
[…]
Proof. W.l.o.g we can assume $j=n$ in the inequality (16). And the inequality (16) is equivalent to
$y\lesseqgtr V\otimes y,$
where $V$ is the $n\times n$-matrix defined by:
$V:=\left(\begin{array}[]{cc}\boldsymbol{I}_{n-1}&\boldsymbol{O}_{n-1,1}\\ D&C\end{array}\right),\;D=\left(\begin{array}[]{ccc}v_{1}&\cdots&v_{n-1}\end{array}\right),\;C=(a^{\otimes(-1)}\otimes v_{n}).$
Now we express $V\otimes y$ as follows:
$V\otimes y=A\otimes y\oplus B\otimes u,$
with :
$A:=\left(\begin{array}[]{cc}\boldsymbol{O}_{n-1,n-1}&\boldsymbol{O}_{n-1,1}\\ \boldsymbol{O}_{1,n-1}&C\end{array}\right),\;B:=\left(\begin{array}[]{c}\boldsymbol{I}_{n-1}\\ D\end{array}\right),\;u=(y_{i})_{i=1}^{n-1}.$
And clearly by Result 3.1 the set $\{y:y\lesseqgtr A\otimes y\oplus B\otimes u\}$ admits a smallest (or greatest) element iff condition (15) is verified. Here, the condition (15) is equivalent to $\lim_{k\rightarrow+\infty}C^{\otimes(k)}=\boldsymbol{O}_{1,1}$ which is equivalent to $a^{\otimes(-1)}\otimes v_{n}<\mathbbm{1}$ because the matrix $C$ has only one elementary circuit $n\rightarrow n$ of weight $a^{\otimes(-1)}\otimes v_{n}$. To conclude we just have to note that by defining $\overline{v}^{\intercal}:=(v_{1},\ldots,v_{n-1},\mbox{{\scriptsize$\mathbb{O}$}})$ the inequality (18a) is verified. $\Box$
4 The minimization problem mpr
At the step 0 of the substitution the cone associated with the constraints system (6c) of the mpr problem is denoted $\mathcal{C}(A^{\;+},A^{\;-})^{[0]}$ and is defined by:
$\mathcal{C}(A^{+},A^{-})^{[0]}:=\{w\in\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}}^{n+2}:A^{+[0]}\otimes w\geq A^{-[0]}\otimes w\}$ (19a)
where:
$(A^{+},A^{-},w)^{[0]}:=\left(\begin{array}[]{ccc}\mathbbm{1}&\boldsymbol{-\infty}^{\intercal}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \boldsymbol{-\infty}&A&b\end{array}\right),\;\left(\begin{array}[]{ccc}\mbox{{\scriptsize$\mathbb{O}$}}&c^{\intercal}&c_{h}\\ \boldsymbol{-\infty}&C&d\end{array}\right),\;\left(\begin{array}[]{c}z\\ x\\ h\end{array}\right).$ (19b)
We use the following conventions.
Numerotation convention 4.1
The rows of the matrices $A^{+},A^{-}$ are numbered from 0 to $m$. The columns of the matrices $A^{+},A^{-}$ are indexed by $j$, $j$ varying from $j=0$ to $j=n+1$. But sometimes it will be useful to use the variables $z,x_{1},\ldots,x_{n},h$. The components of the vector $w$ are numbered from 0 to $n+1$.
The cost function at step 0 of the substitution is denoted and defined by: $\textsf{cost}^{[0]}(x,h):=c^{\intercal}\otimes x\oplus c_{h}\otimes h$, with $c_{h}=\mbox{{\scriptsize$\mathbb{O}$}}$.
The set of the stored linear equalities is denoted $\mathcal{L}^{[k]}$ at each step $k$ of the substitution. And for $k=0$ we have $\mathcal{L}^{[0]}=\emptyset$.
The vector $x$ is called the vector of the remaining variables of the mpr. Of course at step 0 of the method: $x=(x_{j})_{j=1}^{n}$. To $x$ we associate the set of the remaining variables denoted $\mathbbm{x}$. And of course at step 0 of the method: $\mathbbm{x}^{[0]}=\{x_{1},\ldots,x_{n}\}$. We have the following equivalence:
$x_{j}\notin\mathbbm{x}\Leftrightarrow x_{j}=\mbox{{\scriptsize$\mathbb{O}$}}\mbox{ in the vector of remaining variables $x$}.$ (20)
Because $x=\boldsymbol{+\infty}\in\mathcal{P}(A,b,C,d)\cup\{\boldsymbol{\pm\infty}\}$ and because $\min=$-$\infty$ is an acceptable answer for the minimization problem mpr one make the following assumption for the parametrized reseach of a minimum reached by $x$.
Assumption 4.1 (Parametrized reseach domain for the variables $x_{j}$)
For the research of a minimum for $z$ we can assume that $\forall x_{j}\in\mathbbm{x}$: $x_{j}\in[\lambda,+\infty]$ for some arbitrary $\mbox{{\scriptsize$\mathbb{O}$}}\leq\lambda\leq+\infty$.
The homogenization variable $h$ is not substituable (recall). And the general scheme of the substitution for the mpr problem is to arrive after the $n$ substitutions of the variables $x_{j}$ at the following situation:
$\begin{array}[]{l}z\geq c_{h}^{[n]}\otimes h\\ b^{[n]}\otimes h\geq d^{[n]}\otimes h.\end{array}$ (21)
Where $b^{[n]},d^{[n]}$ are two $m$-dimensional column vectors. We then have the following cases.
  • Case $1$: $b^{[n]}\geq d^{[n]}$. In this case $h$ can be set to $\mathbbm{1}$ and $\min(z)=c_{h}^{[n]}$.
  • Case $2$: $b^{[n]}\ngeq d^{[n]}$. In this case $h=+\infty$ is the only possible value for $h$ because $h>\mbox{{\scriptsize$\mathbb{O}$}}$ (see condition (6d) and $\min(z)=+\infty$.
This discussion leads to assume:
Assumption 4.2 (Parametrized domain for $h$)
For the mpr problem we assume that the homogenization variable $h$ has a parametrized interval $\mu,+\infty]$ for some arbitrary $\mbox{{\scriptsize$\mathbb{O}$}}<\mu\leq+\infty$.
In conclusion, the parametrized domain of research of a minimum is de noted $R_{\lambda\mu}(x,h)$ and defined by:
$R_{\lambda\mu}(x,h):=\times_{x_{j}\in\mathbbm{x}}[\lambda,+\infty]\times[\mu,+\infty].$ (22a)
Where because $h$ must be $\mbox{{\scriptsize$\mathbb{O}$}}$ we can assume that the following condition holds:
$\forall\theta,\forall\beta\neq\mbox{{\scriptsize$\mathbb{O}$}},\;\exists\lambda,\mu:\;\theta\otimes\lambda<\beta\otimes\mu.$ (22b)
4.1 Study of the cone $\mathcal{C}(i):=\{z\geq c^{\intercal}\otimes x\oplus c_{h}\otimes h,\;a^{+}_{i,.}\otimes w\geq a^{-}_{i,.}\otimes w\}$
Based on the result of Proposition 3.2 $\forall i=1,\ldots,m$, $\forall j=1,\ldots,n$ such that the following condition is satisfied:
$a^{+}_{i,j}\neq\mbox{{\scriptsize$\mathbb{O}$}}$ and $a^{+}_{i,j}>a^{-}_{i,j}$ (23)
we define the following valid inequality:
$I_{\geq}(x_{j},i):=\{a^{+}_{i,j}\otimes x_{j}\geq a^{-}_{i,.}\otimes w\}.$ (24)
Under condition (23) the inequality $I_{\geq}(x_{j},i)$ can be rewritten as:
$I_{\geq}(x_{j},i)=\{x_{j}\geq f^{\geq}_{ij}(x,h)\}.$ (25)
Where the $(x,h)$-linear function $f^{\geq}_{ij}(x,h)$ is defined by:
$f^{\geq}_{ij}(x,h):=\ell_{ij}(x)\oplus r_{ij}\otimes h,$ (26a)
with $\ell_{ij}$ is the following $x$-linear function which does not depend on $x_{j}$:
$\ell_{ij}(x):=v^{\intercal}_{ij}\otimes x,$ (26b)
where $v_{ij}$ denotes the $n$-dimensional vector such that:
[…]
and the scalar $r_{ij}$ is defined by:
$r_{ij}:=(a^{+}_{i,j})^{\otimes(-1)}\otimes d_{i}.$ (26d)
Assume that $I_{\geq}(x_{j},i)$ is valid. Then, by replacing $x_{j}$ by $f^{\geq}_{ij}(x,h)$ in the cost function $c^{\intercal}\otimes x\oplus c_{h}\otimes h$ we define the following function:
$f_{z,ij}(x,h):=\oplus_{j^{\prime}\neq j}c_{j^{\prime}}\otimes x_{j^{\prime}}\oplus c_{j}\otimes f^{\geq}_{ij}(x,h)\oplus c_{h}\otimes h.$ (27)
By replacing $f^{\geq}_{ij}(x,h)$ in (27) we obtain the new expression for $f_{z,ij}(x,h)$:
$f_{z,ij}(x,h):=\oplus_{j^{\prime}\neq j}(c_{j^{\prime}}\oplus c_{j}\otimes v_{ij,j^{\prime}})\otimes x_{j^{\prime}}\oplus(c_{h}\oplus c_{j}\otimes r_{ij})\otimes h.$ (28)
The next Theorem will be useful to justify the substitution method developed in this paper.
Theorem 4.1 (Saturation of an inequality)
For all $x_{j}$ such that $I_{\geq}(x_{j},i)$ is valid in the sense of Proposition 3.2 we have:
$\inf_{\{x:\;x_{j}\geq f^{\geq}_{ij}(x,h)\}}\left(\begin{array}[]{c}c^{\intercal}\otimes x\oplus c_{h}\otimes h\\ x_{j}\end{array}\right)\;=\left(\begin{array}[]{c}f_{z,ij}(x,h)\\ f^{\geq}_{ij}(x,h)\end{array}\right).$ (29)
And the infimum is reached at $x$ such that we have:
$x_{j}=f^{\geq}_{ij}(x,h).$ (30)
Proof. We have the following implication:
$x_{j}\geq f^{\geq}_{ij}(x,h)\Rightarrow\forall\alpha\in\mathbb{R}_{\mbox{{\scriptsize$\mathbb{O}$}}},\;\alpha\otimes x_{j}\geq\alpha\otimes f^{\geq}_{ij}(x,h).$
In particular, the above implication is true for $\alpha=c_{j}$.
All the $(x,h)$-linear functions are non-decreasing. Noticing that the cost function: $(x,h)\mapsto c^{\intercal}\otimes x\oplus c_{h}\otimes h$ is a $(x,h)$-linear function the result is proved. $\Box$
4.2 How to choose the next variable for substitution in mpr
4.2.1 Classification of the variables and the linear functions of mpr
Based on the result of Proposition 3.2, $\forall i=1,\ldots,m$, $\forall j=1,\ldots,n$ such that the following condition is satisfied:
$a^{-}_{i,j}\neq\mbox{{\scriptsize$\mathbb{O}$}}$ and $a^{+}_{i,j}<a^{-}_{i,j}$ (31)
we define the following valid inequality:
$I_{\leq}(x_{j},i):=\{a^{-}_{i,j}\otimes x_{j}\leq a^{+}_{i,.}\otimes w\}.$ (32)
Under condition (31) the result of Proposition 3.2 allows us to remove $x_{j}$ from the expression $a^{+}_{i,.}\otimes w$. This new $(x,h)$-linear function which does not depend on $x_{j}$ is denoted $f_{ij}^{\leq}(x,h)$. The expression is not given here because the only information to know in the mpr problem is that this function is well defined.
We use the convention that $I_{\geq}(x_{j},i)\neq\emptyset$ (resp. $I_{\leq}(x_{j},i)\neq\emptyset$) equivalently means that $I_{\geq}(x_{j},i)$ (resp. $I_{\leq}(x_{j},i)$) is valid in the sense of Proposition 3.2.
Definition 4.1
A variable $x_{j}$ is said to be
  • $\min$-bounded if $I_{\geq}(x_{j}):=\cup_{i=1}^{m}I_{\geq}(x_{j},i)\neq\emptyset$. Otherwise $x_{j}$ is said to be $\min$-unbounded.
  • $\max$-bounded if $I_{\leq}(x_{j}):=\cup_{i=1}^{m}I_{\leq}(x_{j},i)\neq\emptyset$. Otherwise $x_{j}$ is said to be $\max$-unbounded.
  • dominating if $x_{j}$ is $\min$-and-$\max$-bounded.
And clearly, the dominating variables have a variation domain smaller (in the sense of the set inclusion) than the others variables. So that there exists an order between variables denoted $\preceq_{var}$ based on their variation domain such that:
$\{x_{j}\in\mathbbm{x}:\mbox{$I_{\geq}(x_{j})\neq\emptyset$ and $I_{\leq}(x_{j})\neq\emptyset$}\}\preceq_{var}\{x_{j}\in\mathbbm{x}:\mbox{$I_{\geq}(x_{j})\neq\emptyset$ and $I_{\leq}(x_{j})=\emptyset$}\}.$ (33)
And of course we also mentioned that:
$\{x_{j}\in\mathbbm{x}:\mbox{$I_{\geq}(x_{j})\neq\emptyset$ and $I_{\leq}(x_{j})=\emptyset$}\}\preceq_{var}\{x_{j}\in\mathbbm{x}:\mbox{$I_{\geq}(x_{j})=\emptyset$ and $I_{\leq}(x_{j})=\emptyset$}\}$.
Definition 4.2
A non null $(x,h)$-linear function $f:(x,h)\mapsto\alpha^{\intercal}\otimes x\oplus\beta\otimes h$ is said to be $h$-bounded if $\beta\neq\mbox{{\scriptsize$\mathbb{O}$}}$. Otherwise the function $f$ is said to be $h$-unbounded.
We denote $h\mathbb{B}$ the set of $h$-bounded non null linear functions and we denote $h\mathbb{U}$ the set of non null $h$-unbounded linear functions. And $(h\mathbb{B},h\mathbb{U})$ is a partition of the set of all non null $(x,h)$-linear functions. We have the following interval calculus results.
  • $\forall f=\alpha^{\intercal}\otimes x\oplus\beta\otimes h\in h\mathbb{B}$, using interval calculus formulae we have:
    $f(R_{\lambda\mu}(x,h))=[\oplus_{i}\alpha_{i}\otimes\lambda\oplus\beta\otimes\lambda,+\infty],$ (34a)
    and by assumption ((22b) with $\theta=\oplus_{i}\alpha_{i}$) on $\lambda$ and $\mu$ we have:
    $\forall\alpha\in\mathbb{R}_{\mbox{{\scriptsize$\mathbb{O}$}}}^{n},\forall\beta\neq\mbox{{\scriptsize$\mathbb{O}$}}:\;f(R_{\lambda\mu}(x,h))=[\beta\otimes\mu,+\infty]$ (34b)
    Let us stress that this interval does not depend on the $n$-dimensional row vector $\alpha$ and by assumption (22b):
    $\beta\otimes\mu,+\infty]\subset[\oplus_{i}\alpha_{i}\otimes\lambda,+\infty].$ (34c)
  • And $\forall f=\alpha^{\intercal}\otimes x\in h\mathbb{U}$:
    $f(R_{\lambda\mu}(x,h))=[\oplus_{i}\alpha_{i}\otimes\lambda,+\infty].$ (35)
From this above results we have the following ordering between the non null linear functions denoted $\preceq_{fct}$ which is based on their value domain:
$\{f:f\in h\mathbb{B}\}\preceq_{fct}\{f:f\in h\mathbb{U}\}.$ (36)
4.2.2 Choosing a variable for a new substitution after $k$ substitutions in mpr
In this section we describe the procedure for choosing a remaining variable to be substituted. Assuming that $k$ substitutions have been done the cone associated with the mpr is denoted $\mathcal{C}(A^{+},A^{-})^{[k]}$. The set of linear equalities at step $k$, say $\mathcal{L}^{[k]}$, contains $k$ equalities of the form $\{x_{j}=f\}$ where $f$ is a $(x,h)$-linear function. To lighten notations the superscript $[k]$ will not be applied on the vector of the remaining variables $x$ and the set of the remaining variables $\mathbbm{x}$ in the sequel. Recall that we use numerotation convention 4.1 p. 4.1.
In the next theorem we characterize a normal end of the substitution process.
[…]
The function $h\mapsto c_{h}\otimes h$ is the lower bound of the cost function $(x,h)\mapsto c^{\intercal}_{\mathbbm{x}}\otimes x\oplus c_{h}\otimes h$ which is attained at $x=\boldsymbol{-\infty}$ iff the following condition holds:
$A_{[|1,m|]h}^{+[k]}\geq A_{[|1,m|]h}^{-[k]}.$ (37)
Proof. Considering the system of inequalities $A^{+[k]}\otimes w\geq A^{-[k]}\otimes w$. The $|1,m|]\times\mathbbm{x}\cup\{h\}$ part of the system is expressed as follows:
$A_{[|1,m|]\mathbbm{x}}^{+[k]}\otimes x\oplus A_{[|1,m|]h}^{+[k]}\otimes h\geq A_{[|1,m|]\mathbbm{x}}^{-[k]}\otimes x\oplus A_{[|1,m|]h}^{-[k]}\otimes h.$
We have $z=c_{h}\otimes h$ at $x=\boldsymbol{-\infty}$ iff $(z,\boldsymbol{-\infty},h)\in\mathcal{C}(A^{+},A^{-})^{[k]}$. And based on the above inequality it means that $A_{[|1,m|]h}^{+[k]}\otimes h\geq A_{[|1,m|]h}^{-[k]}\otimes h$. And this last inequality does not imply that $h=+\infty$ is the only solution iff $A_{[|1,m|]h}^{+[k]}\geq A_{[|1,m|]h}^{-[k]}$. And the equivalence is proved. $\Box$
In the next Theorem we present the proof of the optimal choice for one or several variables which can be substituted. Each variable is associated with an inequality which belongs to a given set of inequalities.
Theorem 4.3
Let $\mathbb{I}$ be a subset of valid inequalities $I_{\geq}(x_{j},i)$, ie. a set such that:
$\mathbb{I}\subseteq I_{\geq}(\mathbbm{x}).$ (38)
Where
$I_{\geq}(\mathbbm{x}):=\{(i,j):\mbox{ $x_{j}\in\mathbbm{x}$ and $I_{\geq}(x_{j},i)$ is valid}\}.$ (39)
We have the following implication:
 $\mbox{ $I_{\geq}(x_{j},i)\neq\emptyset$ or equivalently $I_{\geq}(x_{j},i)$ is valid}\Rightarrow z\geq f_{z,ij}(x,h).$ (40)
Recalling that $f_{z,ij}(x,h)$ is defined by (28).
Let $\mathbb{T}$ and $\mathbb{T}^{\prime}$ denoting either $\mathbb{B}$ or $\mathbb{U}$, respectively. The $2$-tuple $(\mathbb{T},\mathbb{T}^{\prime})$ is defined as a function of the set $\mathbb{I}$ according to $\preceq_{fct}$ as follows (recalling the priority to the cost function):
[…]
Then, define the following set of indexes:
$\mathcal{T}(\mathbb{T},\mathbb{T}^{\prime},\mathbb{I}):=\{(i,j)\in\mathbb{I}:(f_{z,ij},f^{\geq}_{ij})\in(h\mathbb{T},h\mathbb{T}^{\prime})\}.$ (42)
Let us define the following assertion which deals with the cost $z$ and an interval $\tau$:
$\begin{array}[]{l}\mathrm{Z}(\tau):\\ \forall(i,j)\in\mathcal{T}(\mathbb{T},\mathbb{T}^{\prime},\mathbb{I})\;\forall z\;(z\in\tau\Rightarrow\exists(x,h)\in R_{\lambda\mu}(x,h)\;z\geq f_{z,ij}(x,h)).\end{array}$ (43)
And let us define the following assertion dealing with variables of the problem and intervals $\tau,\tau^{\prime}$:
$\begin{array}[]{l}\mathrm{X}(\tau,\tau^{\prime}):\\ \forall(i,j)\in\mathcal{T}(\mathbb{T},\mathbb{T}^{\prime},\mathbb{I})\cap\textsf{argMin}(\tau)\;\forall x_{j}\;(x_{j}\in\tau^{\prime}\Rightarrow\exists(x,h)\in R_{\lambda\mu}(x,h)\;x_{j}\geq f^{\geq}_{ij}(x,h)).\end{array}$ (44)
Then, we have the following equivalences.
$\mathrm{Z}(\tau)\mbox{ true }\Leftrightarrow\tau=\textsf{Min}\{f_{z,ij}(R_{\lambda\mu}(x,h)),(i,j)\in\mathcal{T}(\mathbb{T},\mathbb{T}^{\prime},\mathbb{I})\}.$ (45)
And
$\mathrm{X}(\tau,\tau^{\prime})\mbox{ true }\Leftrightarrow\tau^{\prime}=\textsf{Min}\{f^{\geq}_{ij}(R_{\lambda\mu}(x,h)),(i,j)\in\textsf{argMin}(\tau)\cap\mathcal{T}(\mathbb{T},\mathbb{T}^{\prime},\mathbb{I})\}.$ (46)
Proof. The proof of the implication (40) is a straightforward consequence of Theorem 4.1.
The equivalence (45) comes from the properties of the familly of intervals we are considering in this paper. We have a total order (see Proposition 3.1) and the Min of a set of intervals coincides with the intersection of the same set of intervals (see (13)). The proof is achieved by noticing that $\forall(i,j)\in\textsf{argMin}(\tau)\cap\mathcal{T}(\mathbb{T},\mathbb{T}^{\prime},\mathbb{I})$ the variables $x_{j}$ can be exchanged. And thus we have to take the Min. $\Box$
We are now in position to indicate how to choose a remaining variable to be substituted. The choice also depends on the following different cases which are listed hereafter.
Case 0: $min(z)=+\infty$ or no minimum. Case $0.1$: $\mathbbm{x}=\emptyset$ and the condition (37) of Theorem 4.2 does not holds. $\mathbbm{x}=\emptyset$ means that $k=n$ (no more variable to be substituted) and $\textsf{cost}^{[n]}(x,h)=c_{h}^{[n]}\otimes h$. The condition (37) of Theorem 4.2 does not holds means $A_{[|1,m|]h}^{+[n]}\ngeq A_{[|1,m|]h}^{-[n]}$ then $h=+\infty$ is the only possible value for $h$ and by backward substitution we have $z=+\infty$, $x=\boldsymbol{+\infty}$.
Case $0.2$: $\mathbbm{x}\neq\emptyset$ and $I_{\leq}(\mathbbm{x})=\emptyset$ and $I_{\geq}(\mathbbm{x})=\emptyset$ and the condition (37) of Theorem 4.2 does not holds. The substitution process stops because no valid inequalities can be generated. No minimum reached.
Case $1$: switching case, cf. § 2.4. $\mathbbm{x}^{+}=\emptyset$ and $\textsf{cost}^{[k]}(x,h)=c_{h}^{[k]}\otimes h$ and the condition (37) of Theorem 4.2 does not holds and $\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}\neq\emptyset$ and $I_{\geq}(\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}})=\emptyset$ and $I_{\leq}(\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}})\neq\emptyset$. The substitution must switch to the following maximizing problem:
$\textsf{PMR}(A,b,C,d,c,c_{h},z,x,h,(m,n)),$ (47a)
where the matrices $A$ and $C$ are defined by:
$A:=A^{-[k]}_{[|1,m|][|1,n|]},\;C:=A^{+[k]}_{[|1,m|][|1,n|]}.$ (47b)
The vectors $b$ and $d$ are defined by:
$b:=A^{-[k]}_{[|1,m|]h},d:=A^{+[k]}_{[|1,m|]h}$. (47c)
The initial cost function of the MPR is characterized by the vector $c$ and the constant $c_{h}$ such that:
$c:=\boldsymbol{-\infty},\;c_{h}:=c_{h}^{[k]}.$ (47d)
And the variables of the problem are:
$z:=z,\;x\leftarrow x^{\mbox{{\scriptsize$\mathbb{O}$}}},\;h:=h.$ (47e)
Where the notation $x\leftarrow x^{\mbox{{\scriptsize$\mathbb{O}$}}}$ means that we store in the $n$-dimensional vector $x$ the remaining variables of the set $\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}$. The other components (already substituted variables) are set to $\mathbb{O}$.
And the dimensions $(m,n)$ of the MPR are obviously the same as in the mpr.
Case $2$: Not Case 0 and not Case $1$. The first case is Case $2.1$: $\mathbbm{x}=\emptyset$ and the condition (37) of Theorem 4.2 holds. This means that $k=n$ (no more variable to be substituted) and $\textsf{cost}^{[n]}(x,h)=c_{h}^{[n]}\otimes h$ and $A_{[|1,m|]h}^{+[n]}\geq A_{[|1,m|]h}^{-[n]}$ then $h$ can take all possible values in particular we can take $h=\mathbbm{1}$ and the mpr succeeds with finite minimum $z=c_{h}^{[n]}$.
The other case is Case $2.2$: $\mathbbm{x}\neq\emptyset$ and $\exists x_{j}\in\mathbbm{x}$ such that: $I_{\geq}(x_{j})\neq\emptyset$. Let us distinguish the following subcases which are based on the form of the cost function:
$\textsf{cost}^{[k]}(x,h)=c^{\intercal}_{\mathbbm{x}^{+}}\otimes x^{+}\oplus c^{\intercal}_{\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}}\otimes x^{\mbox{{\scriptsize$\mathbb{O}$}}}\oplus c_{h}\otimes h,$ (48a)
where:
$\mbox{$\mathbbm{x}^{+}:=\{x_{j}\in\mathbbm{x}:\;c_{j}>\mbox{{\scriptsize$\mathbb{O}$}}\}$ and $\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}:=\{x_{j}\in\mathbbm{x}:\;c_{j}=\mbox{{\scriptsize$\mathbb{O}$}}\}$}.$ (48b)
And the vectors $x^{+}$ and $x^{\mbox{{\scriptsize$\mathbb{O}$}}}$ are defined by
$x^{+}:=(x_{j})_{\{j:x_{j}\in\mathbbm{x}^{+}\}}\mbox{ and }x^{\mbox{{\scriptsize$\mathbb{O}$}}}:=(x_{j})_{\{j:x_{j}\in\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}\}}.$ (48c)
[…]
Case $2.2.2$: otherwise, the condition (37) of Theorem 4.2 does not hold. Based on the hierarchy $\preceq_{var}$ (see (33)) between the remaining variables let us define the set of the dominating variables $D^{+\mbox{{\scriptsize$\mathbb{O}$}}}:=\{x_{j}\in\mathbbm{x}:\;I_{\leq}(x_{j})\neq\emptyset\mbox{ and }I_{\geq}(x_{j})\neq\emptyset\}$. And the set of the valid inequalities:
[…]
Where: $I_{\geq}(\mathbbm{x}\cap D^{+\mbox{{\scriptsize$\mathbb{O}$}}}):=\{(i,j):\mbox{ $x_{j}\in\mathbbm{x}\cap D^{+\mbox{{\scriptsize$\mathbb{O}$}}}$ and $I_{\geq}(x_{j},i)$ is valid}\}$.
We apply Theorem 4.3 recalling that the $2$-tuple $(\mathbb{T},\mathbb{T}^{\prime})$ is defined according with $\preceq_{fct}$ by (41) and the subset of valid inequalities is $\mathbb{I}:=\mathbb{I}^{+\mbox{{\scriptsize$\mathbb{O}$}}}(D)$. Thus, the “working space” of indexes of rows and columns (or variables) of the system of inequalities is: $\mathcal{T}(\mathbb{T},\mathbb{T}^{\prime},\mathbb{I}^{+\mbox{{\scriptsize$\mathbb{O}$}}}(D))$.
Thus, we compute the interval $\tau$ such that $\mathrm{Z}(\tau)$ is true (see (43) and (45)).
If $\textsf{argMin}(\tau)$ is a singleton $\{(i^{*},j^{*})\}$ then take the unique $2$-tuple $(i^{*},j^{*})$ for the substitution of $x_{j^{*}}=f^{\geq}_{i^{*}j^{*}}(x,h)$. Otherwise, take any $2$-tuple $(i^{*},j^{*})$ element of $\textsf{argMin}(\tau^{\prime})$ recalling that the interval $\tau^{\prime}$ is the interval such that $\mathrm{X}(\tau,\tau^{\prime})$ is true (see (44) and (46)).
The new set of equalities $\mathcal{L}^{[k+1]}$, the new cost function $\text{cost}^{[k+1]}(x,h)$ and the new cone $\mathcal{C}(A^{+},A^{-})^{[k+1]}$ are defined by the procedure described in section 4.2.3.
4.2.3 The new characteristic elements of the mpr after substitution
The new set of equalities is defined as follows. If conditions of Theorem 4.2 are verified then set $k+1=n$ and $\mathcal{L}^{[n]}:=\mathcal{L}^{[k]}\cup_{x_{j}\in\mathbbm{x}}\{x_{j}=\mbox{{\scriptsize$\mathbb{O}$}}\}\cup\{z=c_{h}\otimes h\}$. Then, pmrp stops and we solve the linear system $\mathcal{L}^{[n]}$ and express all the $x_{j}$’s as function of the homogenization variable $h$. Otherwise, we define $\mathcal{L}^{[k+1]}$ as: $\mathcal{L}^{[k+1]}:=\mathcal{L}^{[k]}\cup\{x_{j^{*}}=f_{i^{*}j^{*}}(x,h)\}$ and we apply Theorem 4.1 of § 4.1. And based on the previous cases the new cost function is defined as:
$\textsf{cost}^{[k+1]}(x,h):=f_{z,i^{*}j^{*}}(x,h).$ (50)
And in all cases the new cone $\mathcal{C}(A^{+},A^{-})^{[k+1]}$ associated with pmrp is deduced from the following set of inequalities:
[…]
and $\ell$ is the $n+2$-dimensional column vector which has the same components as the vector $w^{[k]}$ except its $j^{*}$th component wich is:
$\ell_{j^{*}}:=f_{i^{*}j^{*}}(x,h).$ (51b)
Let us define the following $(n+2)\times(n+2)$ transition matrix $T^{k\rightarrow k+1}$ by:
$\forall j:t^{k\rightarrow k+1}_{j,.}:=\left\{\begin{array}[]{lc}e^{\intercal}_{j}\mbox{ if $j\neq j^{*}$}\\ (\mbox{{\scriptsize$\mathbb{O}$}},v^{\intercal}_{i^{*}j^{*}},r_{i^{*}j^{*}})&\mbox{ if $j=j^{*}$.}\end{array}\right$. (52)
Where $e^{\intercal}_{j}$ denotes the $j$th $n+2$-dimensional row vector of the $(n+2)\times(n+2)$-identity matrix $\boldsymbol{I}_{n+2}$.
The new matrices $(A^{+},A^{-})^{[k+1]}$ are defined by:
$(A^{+},A^{-})^{[k+1]}:=\textsf{setrowtozero}(A^{+[k]}\otimes T^{k\rightarrow k+1},A^{-[k]}\otimes T^{k\rightarrow k+1})$ (53)
recalling that function setrowtozero is defined by (7).
And the new vector $w^{[k+1]}$ is defined by:
[…]
And the new set $\mathbbm{x}$ of the remaining variables is defined by:
$\mathbbm{x}:=\mathbbm{x}\setminus\{x_{j^{*}}\}.$ (55)
5 Linear fractional programming
In this section we consider the following $(\max,+)$-analogue of the linear fractional programming (see [6]) defined as follows.
$\min\{z=p^{\intercal}\otimes x/r^{\intercal}\otimes x,x\in\mathcal{P}(A,b,C,d)\cup\{\boldsymbol{\pm\infty}\}\}.$ (56a)
Where $p$, $r$ are two $n$-dimensional vectors satisfying the following assumption:
$\mbox{{\bf Assumption $\boldsymbol{1^{f}}$}}:\;\mbox{$p\neq\boldsymbol{-\infty}$ and $r\neq\boldsymbol{-\infty}$}.$ (56b)
This kind of problem has been studied in eg. [10] and [11] where pseudo-polynomial algorithms have been provided.
Hereafter, we follow the $(\max,+)$-analogue of the Charnes-Cooper transformation [6] which has been used in [11]. Using also the Fourier’s trick the problem (56a) is unfolded as follows:
[…]
Remark 5.1
In [11] the authors follow exactly the Charnes-Cooper transformation. This means that the $(y)$-polyhedron part is not an inequality but the equality: $r^{\intercal}\otimes y=\mathbbm{1}$. This latter equality is equivalent to: (1). $r^{\intercal}\otimes y\geq\mathbbm{1}$ and (2). $\mathbbm{1}\geq r^{\intercal}\otimes y$. But substitution allows us to drop the inequality (2) because when an inequality is chosen using our criteria developed in § 4.2.2 then we use the result of Theorem 4.1 of § 4.1 to saturate this inequality, ie. $\geq$ becomes ==.
Now, we just have to homogenize the $(y)$-polyhedron part by adding the homogeneous variable $h$ and we obtain the following mpr problem:
$\begin{array}[]{l}\min(z)\\ z\geq p^{\intercal}\otimes y\;:\mbox{Fourier's trick}.\end{array}$ (58a)
The research of a minimum is developed on the following cone:
$\begin{array}[]{l}A\otimes y\oplus b\otimes t\geq C\otimes y\oplus d\otimes t\;:\mbox{$(y,t)$-cone part}\\ r^{\intercal}\otimes y\geq h\;:\mbox{$(y)$-polyhedron part homogenized}.\end{array}$ (58b)
The variable $t$ used in the Charnes-Cooper transformation is such that:
$t\geq\mbox{{\scriptsize$\mathbb{O}$}},\;y=t\otimes x.$ (58c)
And the homogenization variable $h$ satifies:
$h>\mbox{{\scriptsize$\mathbb{O}$}}.$ (58d)
And
$\mbox{$z$ and $h$ are not substituable}.$ (58e)
Finally, among the variables of the problem $z,y_{1},\ldots,y_{n},t$ we have:
[…]
6 Numerical examples
In ths section we study pedagogical examples. In § 6.1 we illustrate the coherency of our strongly polynomial method on a toy example.
In § 6.2 (problem of minimization) and in § 6.3 (problem of fractional minimization) we illustrate our strongly polynomial method on some numerical examples borrowed from the litterature which are solved using pseudo-polynomial method in [10].
6.1 A minimum which is equal to $\infty$
In this subsection we consider the following minimization problem: the homogenization variable $h$:
$\min\{z=x_{1},x=(x_{1})\in\mathcal{P}(A,b,C,d)\cup\{\boldsymbol{\pm\infty}\}\},$
where the polyhedron $\mathcal{P}(A,b,C,d)$ is characterized by the following vectors and matrices:
$A:=\left(\begin{array}[]{c}\mathbbm{1}\\ 2^{\otimes(-1)}\end{array}\right),\;b:=\left(\begin{array}[]{c}\mathbbm{1}\\ 1\end{array}\right),\;C:=\left(\begin{array}[]{c}1^{\otimes(-1)}\\ \mathbbm{1}\end{array}\right),\;d:=\left(\begin{array}[]{c}2\\ \mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right).$
Because the example is very small we do not need to describe the cone associated with mpr. Using the homogenization variable $h$ the mpr is unfolded as follows:
$\begin{array}[]{l}\min(z)\\ l_{0}:\;z\geq x_{1}\\ l_{1}:\;x_{1}\oplus h\geq 1^{\otimes(-1)}\otimes x_{1}\oplus 2\otimes h\\ l_{2}:\;2^{\otimes(-1)}\otimes x_{1}\oplus 1\otimes h\geq x_{1}.\end{array}$ (59)
Only $l_{1}$ generates the valid inequality $I_{\geq}(x_{1},1)=\{x_{1}\geq 2\otimes h\}$. Thus, the substitution of $x_{1}$ by $2\otimes h$ (ie. we have $\mathcal{L}^{[1]}=\{x_{1}=2\otimes h\}$) provides the new system:
$\begin{array}[]{l}\min(z)\\ l_{0}:\;z\geq 2\otimes h\\ l_{1}:\;2\otimes h\geq 2\otimes h\\ l_{2}:\;1\otimes h\geq 2\otimes h\end{array}$ (60)
Constraint $l_{2}$ implies that $h=+\infty$ is the only possible value for $h$, where $h$ is assumed to be $\mbox{{\scriptsize$\mathbb{O}$}}$. Then, by backward substitution we have: $z=2\otimes+\infty=+\infty$ and $x_{1}=2\otimes+\infty=+\infty$. And $\infty$ is element of the set $\mathcal{P}(A,b,C,d)\cup\{\boldsymbol{\pm\infty}\}$.
This illustrates the coherency of the method because $\min(z)=+\infty$ is actually reached at $x=(+\infty)$ which is a point of the polyhedron $\cup\{\boldsymbol{\pm\infty}\}$.
6.2 Minimization Example
In this subsection we consider the Example 2 p. 1463 of [10] which is:
$\min\{z=2\otimes x_{1}\oplus 4^{\otimes(-1)}\otimes x_{2},x=(x_{1},x_{2})^{\intercal}\in\mathcal{P}(A,b,C,d)\cup\{\boldsymbol{\pm\infty}\}\},$
Where:
$A:=\left(\begin{array}[]{cc}2^{\otimes(-1)}&\mathbbm{1}\\ \mathbbm{1}&1^{\otimes(-1)}\\ 1&2^{\otimes(-1)}\\ 2&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ 2^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}\\ 4^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right),\;b:=\left(\begin{array}[]{c}\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}\\ \mathbbm{1}\\ \mathbbm{1}\\ \mathbbm{1}\end{array}\right),\;C:=\left(\begin{array}[]{cc}\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&4^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&5^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&6^{\otimes(-1)}\end{array}\right),\;d:=\left(\begin{array}[]{c}\mathbbm{1}\\ \mathbbm{1}\\ \mathbbm{1}\\ \mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right).$
with the following change of notations $A\leftrightarrow B$, $b\leftrightarrow d$, $C\leftrightarrow A$ and $d\leftrightarrow c$.
The mpr is defined on the cone $\mathcal{C}(A,b,C,d)$ obtained from the polyhedron $\mathcal{P}(A,b,C,d)$ by adding the homogenization variable $h$ as already explained previuously.
At step 0 of the substitution the cone associated with the above constraints of the mpr problem is denoted $\mathcal{C}(A^{\;+},A^{\;-})^{[0]}$ and is defined by the following system of inequalities:
$A^{+[0]}\otimes w\geq A^{+[0]}\otimes w,$ (61a)
with $w=(z\;x_{1}\;x_{2}\;h)\in\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}}^{4}$ and the matrices $(A^{\;+},A^{\;-})^{[0]}$ are defined by:
$A^{+[0]}:=\left(\begin{array}[]{cccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&2^{\otimes(-1)}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&1^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&1&2^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&2&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&2^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&4^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\end{array}\right),\;A^{-[0]}:=\left(\begin{array}[]{cccc}\mbox{{\scriptsize$\mathbb{O}$}}&2&4^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&4^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&5^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&6^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right).$ (61b)
And the Fourier’s trick for this example is:
$\min(z)$ (62a)
$z\geq\textsf{cost}^{[0]}(x,h)$ (62b)
$\textsf{cost}^{[0]}(x,h)=2\otimes x_{1}\oplus 4^{\otimes(-1)}\otimes x_{2}\oplus\mbox{{\scriptsize$\mathbb{O}$}}\otimes h.$ (62c)
$\mathbbm{x}=\{x_{1},x_{2}\}\neq\emptyset$ and the mpr problem is bounded because one can check that: $\forall x_{j}\in\mathbbm{x}$: $I_{\geq}(x_{j})\cup I_{\leq}(x_{j})\neq\emptyset$. We have $\mathbbm{x}^{+}=\{x_{1},x_{2}\}$ and $\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}=\emptyset$. The set of dominating variables is $D^{+\mbox{{\scriptsize$\mathbb{O}$}}}=\{x_{2}\}$. Indeed, it is easy to see here that $I_{\leq}(x_{1})=\emptyset$ because $A_{[|1,m|]x_{1}}^{-[0]}=\boldsymbol{-\infty}$. Clearly, the condition (37) of Theorem 4.2 does not hold and Case $2.2.2$ applies. We have (see 49): $\mathbb{I}^{+\mbox{{\scriptsize$\mathbb{O}$}}}(D)=I_{\geq}(x_{2})=\{(1,2),(2,2),(3,2)\}$.
And we have the following three possibilities for the substitution of the dominating variable $x_{2}$ in the system of inequalities detailed hereafter:
$\begin{array}[]{lllllll}\left(\begin{array}[]{c}z\\ x_{2}\end{array}\right)_{0,1}&\geq&\left(\begin{array}[]{c}2\otimes x_{1}\oplus 4^{\otimes(-1)}\otimes h\\ h\end{array}\right)&::&\left(\begin{array}[]{c}\mbox{$[4^{\otimes(-1)}\otimes\mu,+\infty]$}\\ \mbox{$[\mathbbm{1}\otimes\mu,+\infty]$}\end{array}\right)&::&\left(\begin{array}[]{c}h\mathbb{B}\\ h\mathbb{B}\end{array}\right)\\ \left(\begin{array}[]{c}z\\ x_{2}\end{array}\right)_{0,2}&\geq&\left(\begin{array}[]{c}2\otimes x_{1}\oplus 3^{\otimes(-1)}\otimes h\\ 1\otimes h\end{array}\right)&::&\left(\begin{array}[]{c}\mbox{$[3^{\otimes(-1)}\otimes\mu,+\infty]$}\\ \mbox{$[1\otimes\mu,+\infty]$}\end{array}\right)&::&\left(\begin{array}[]{c}h\mathbb{B}\\ h\mathbb{B}\end{array}\right)\\ \left(\begin{array}[]{c}z\\ x_{2}\end{array}\right)_{0,3}&\geq&\left(\begin{array}[]{c}2\otimes x_{1}\oplus 2^{\otimes(-1)}\otimes h\\ 2\otimes h\end{array}\right)&::&\left(\begin{array}[]{c}\mbox{$[2^{\otimes(-1)}\otimes\mu,+\infty]$}\\ \mbox{$[2\otimes\mu,+\infty]$}\end{array}\right)&::&\left(\begin{array}[]{c}h\mathbb{B}\\ h\mathbb{B}\end{array}\right).\end{array}$
Applying Theorem 4.3 where from the above array we have $(\mathbb{T},\mathbb{T}^{\prime})=(\mathbb{B},\mathbb{B})$ the minimum interval $\tau$ is $\tau=[2^{\otimes(-1)}\otimes\mu,+\infty]$. And $\textsf{argMin}(\tau)=\{(3,2)\}$ which is a singleton. Thus, $(i^{*},j^{*})=(3,2)$ which means that $x_{2}=2\otimes h$ in row $3$ of the linear system of inequalities (61a). We apply the results of § 4.2.3. Then, the set of linear equalities is now:
$\mathcal{L}^{[1]}=\{x_{2}=2\otimes h\}.$ (63)
The new cost function is:
$\textsf{cost}^{[1]}(x,h):=f_{z,32}(x,h)=2\otimes x_{1}\oplus 2^{\otimes(-1)}\otimes h.$ (64)
The $4\times 4$-transition matrix $T^{0\rightarrow 1}$ is defined by:
$T^{0\rightarrow 1}:=\left(\begin{array}[]{cccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&2\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\end{array}\right).$ (65)
We calculate the following matrices $A^{+}:=A^{+[0]}\otimes T^{0\rightarrow 1}$ and $A^{-}:=A^{-[0]}\otimes T^{0\rightarrow 1}$ and we obtain:
$A^{+}:=\left(\begin{array}[]{cccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&2^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&2\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&1\\ \mbox{{\scriptsize$\mathbb{O}$}}&1&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&2&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&2^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&4^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\end{array}\right),\;A^{-}:=\left(\begin{array}[]{cccc}\mbox{{\scriptsize$\mathbb{O}$}}&2&\mbox{{\scriptsize$\mathbb{O}$}}&2^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&2^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&4^{\otimes(-1)}\end{array}\right).$
We remark that $\forall i\neq 0,4$ we have $a^{+}_{i,.}\geq a^{-}_{i,.}$ thus the new cone $\mathcal{C}(A^{\;+},A^{\;-})^{[1]}$ is defined by:
$A^{+[1]}\otimes w\geq A^{+[1]}\otimes w,$ (66a)
with $w=(z,x_{1},\mbox{{\scriptsize$\mathbb{O}$}},h)^{\intercal}\in\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}}^{4}$ and the matrices $(A^{\;+},A^{\;-})^{[1]}$ are defined as the result of $\textsf{setrowtozero}(A^{+},A^{-})$:
$A^{+[1]}:=\left(\begin{array}[]{cccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&2&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right),\;A^{-[1]}:=\left(\begin{array}[]{cccc}\mbox{{\scriptsize$\mathbb{O}$}}&2&\mbox{{\scriptsize$\mathbb{O}$}}&2^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right).$ (66b)
$\mathbbm{x}=\{x_{1}\}\neq\emptyset$ and the mpr problem is bounded because one can check eg. that: $I_{\geq}(x_{1})\neq\emptyset$. We have $\mathbbm{x}^{+}=\{x_{1}\}$ and of course $\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}=\emptyset$. The set of dominating variables is $D^{+\mbox{{\scriptsize$\mathbb{O}$}}}=\emptyset$. The condition (37) of Theorem 4.2 does not hold and Case $2.2.2$ applies. We have (see 49): $\mathbb{I}^{+\mbox{{\scriptsize$\mathbb{O}$}}}(D)=I_{\geq}(x_{2})=\{(4,1)\}$.
Thus, obviously $(i^{*},j^{*})=(4,1)$ which means that $x_{1}=2^{\otimes(-1)}\otimes h$, in the row $4$ of the linear system of inequalities (66b). We apply the results of § 4.2.3. Then, the set of linear equalities is now:
$\mathcal{L}^{[2]}=\{x_{2}=2\otimes h,x_{1}=2^{\otimes(-1)}\otimes h\}.$ (67)
The new cost function is:
$\textsf{cost}^{[2]}(x,h):=f_{z,32}(x,h)=h.$ (68)
The $4\times 4$-transition matrix $T^{1\rightarrow 2}$ is defined by:
$T^{1\rightarrow 2}:=\left(\begin{array}[]{cccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&2^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\end{array}\right).$ (69)
We calculate the following matrices $A^{+}:=A^{+[1]}\otimes T^{1\rightarrow 2}$ and $A^{-}:=A^{-[1]}\otimes T^{1\rightarrow 2}$ and we obtain:
$A^{+}:=\left(\begin{array}[]{cccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right),\;A^{-}:=\left(\begin{array}[]{cccc}\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right).$
with $w=(z,\mbox{{\scriptsize$\mathbb{O}$}},\mbox{{\scriptsize$\mathbb{O}$}},h)^{\intercal}\in\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}}^{4}$ and the matrices $(A^{\;+},A^{\;-})^{[2]}$ are defined as a result of $\textsf{setrowtozero}(A^{+},A^{-})$:
$A^{+[2]}:=\left(\begin{array}[]{cccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right),\;A^{-[2]}:=\left(\begin{array}[]{cccc}\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right).$ (70)
And here we have: $\boldsymbol{-\infty}=A_{[|1,7|]h}^{+[2]}\geq A_{[|1,7|]h}^{-[2]}=\boldsymbol{-\infty}$. Thus, the reachability condition (37) of Theorem 4.2 is verified. So, mpr has the following solution:
$z=h,\;x_{1}=2^{\otimes(-1)}\otimes h,\;x_{2}=2\otimes h.$
And we retrieve the solution of the Example 2 p. 1463 of [10].
6.3 Fractional Example
In this subsection we illustrate and compare our strongly polynomial method with the pseudo-polynomial method applied on Example 3 p. 1469 of [10]. The starting point of the pseudo-polynomial method is not a polyhedron but the homogenized cone of the following set $\mathcal{P}(A,b,C,d)\cup\{\boldsymbol{\pm\infty}\}$ defined by:
$A\otimes x\oplus b\geq C\otimes x\oplus d,$ (71)
with
$A:=\left(\begin{array}[]{ccc}\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right),\;b:=\left(\begin{array}[]{c}\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}\\ 3\end{array}\right),\;C:=\left(\begin{array}[]{ccc}3^{\otimes(-1)}&4^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}\\ 1^{\otimes(-1)}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ 1&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\end{array}\right),\;d:=\left(\begin{array}[]{c}\mbox{{\scriptsize$\mathbb{O}$}}\\ 1\\ \mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right).$ (72)
With the following change of notations $(A\;b)\leftrightarrow D$ and $(C\;d)\leftrightarrow C$.
The fractional minimization problem fmpr we solve here is defined as:
$\min\{z=x_{2}/3\otimes x_{1},\;x=(x_{1},x_{2},x_{3})^{\intercal}\in\mathcal{P}(A,b,C,d)\}.$ (73)
Applying the results of § 5, the previous fmpr is transformed into the mpr problem defined by (58a)-(58b)-(58c)-(58d) which is specified hereafter.
$\begin{array}[]{l}\min(z)\\ z\geq\textsf{cost}^{[0]}(y,t,h):=y_{2}\end{array}$ (74)
$w:=(z,y_{1},y_{2},y_{3},t,h)^{\intercal}\in\mathcal{C}(A^{+},A^{-})^{[0]}$ (75)
$(A^{+},A^{-})^{[0]}:=\left(\begin{array}[]{cccccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&3&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right),\;\left(\begin{array}[]{cccccc}\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}&4^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&1^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&1&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&1&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\end{array}\right).$ (76)
The set $\{y_{1},y_{2},y_{3},t\}$ is abbreviated by $(\mathbbm{y},t)$.
The mpr is bounded with $I_{\geq}(y_{3})=\emptyset$. $(\mathbbm{y},t)^{+}=\{y_{2}\}$ and $(\mathbbm{y},t)^{\mbox{{\scriptsize$\mathbb{O}$}}}=\{y_{1},y_{3},t\}$. The dominating variables are in $D^{+\mbox{{\scriptsize$\mathbb{O}$}}}=\{y_{1},y_{2},t\}$. The condition (37) of Theorem 4.2 does not hold and Case $2.2.2$ applies. We have (see 49): $\mathbb{I}^{+\mbox{{\scriptsize$\mathbb{O}$}}}(D)=\{(1,t),(2,2),(3,1),(4,t),(5,1)\}$. We enumerate all the possibilities in the next array.
$\begin{array}[]{lllllll}\left(\begin{array}[]{c}z\\ t\end{array}\right)_{0,1}&\geq&\left(\begin{array}[]{c}y_{2}\\ 3^{\otimes(-1)}\otimes y_{1}\oplus 4^{\otimes(-1)}\otimes y_{2}\end{array}\right)&::&\left(\begin{array}[]{c}\mbox{$[\mathbbm{1}\otimes\lambda,+\infty]$}\\ \mbox{$[3^{\otimes(-1)}\otimes\lambda,+\infty]$}\end{array}\right)&::&\left(\begin{array}[]{c}h\mathbb{U}\\ h\mathbb{U}\end{array}\right)\\ \left(\begin{array}[]{c}z\\ y_{2}\end{array}\right)_{0,2}&\geq&\left(\begin{array}[]{c}1^{\otimes(-1)}\otimes y_{1}\oplus 1\otimes t\\ 1^{\otimes(-1)}\otimes y_{1}\oplus 1\otimes t\end{array}\right)&::&\left(\begin{array}[]{c}\mbox{$[1\otimes\lambda,+\infty]$}\\ \mbox{$[1\otimes\lambda,+\infty]$}\end{array}\right)&::&\left(\begin{array}[]{c}h\mathbb{U}\\ h\mathbb{U}\end{array}\right)\\ \left(\begin{array}[]{c}z\\ y_{1}\end{array}\right)_{0,3}&\geq&\left(\begin{array}[]{c}y_{2}\\ t\end{array}\right)&::&\left(\begin{array}[]{c}\mbox{$[\mathbbm{1}\otimes\lambda,+\infty]$}\\ \mbox{$[\mathbbm{1}\otimes\lambda,+\infty]$}\end{array}\right)&::&\left(\begin{array}[]{c}h\mathbb{U}\\ h\mathbb{U}\end{array}\right)\\ \left(\begin{array}[]{c}z\\ t\end{array}\right)_{0,4}&\geq&\left(\begin{array}[]{c}y_{2}\\ 2^{\otimes(-1)}\otimes y_{1}\oplus 3^{\otimes(-1)}\otimes y_{3}\end{array}\right)&::&\left(\begin{array}[]{c}\mbox{$[\mathbbm{1}\otimes\lambda,+\infty]$}\\ \mbox{$[2^{\otimes(-1)}\otimes\lambda,+\infty]$}\end{array}\right)&::&\left(\begin{array}[]{c}h\mathbb{U}\\ h\mathbb{U}\end{array}\right)\\ \left(\begin{array}[]{c}z\\ y_{1}\end{array}\right)_{0,5}&\geq&\left(\begin{array}[]{c}y_{2}\\ 3^{\otimes(-1)}\otimes h\end{array}\right)&::&\left(\begin{array}[]{c}\mbox{$[\mathbbm{1}\otimes\lambda,+\infty]$}\\ \mbox{$[3^{\otimes(-1)}\otimes\mu,+\infty]$}\end{array}\right)&::&\left(\begin{array}[]{c}h\mathbb{U}\\ h\mathbb{B}\end{array}\right)\end{array}$
By Theorem 4.3 we have: $(\mathbb{T},\mathbb{T}^{\prime})=(\mathbb{U},\mathbb{B})$, thus: $(i^{*},j^{*})=(5,1)$. This corresponds to $\left(\begin{array}[]{c}z\\ y_{1}\end{array}\right)_{0,5}$ in the array just above.
The new cost function is: $\textsf{cost}^{[1]}(y,t,h)=y_{2}$. And we define the $6\times 6$-transition matrix $T^{0\rightarrow 1}$ as follows:
$T^{0\rightarrow 1}:=\left(\begin{array}[]{cccccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\end{array}\right).$
The set of linear equalities $\mathcal{L}^{[1]}$ is:
$\mathcal{L}^{[1]}=\mathcal{L}^{[0]}\cup\{y_{1}=3^{\otimes(-1)}\otimes h\}=\{y_{1}=3^{\otimes(-1)}\otimes h\}.$ (77)
We calculate the matrices $A^{+}=A^{+[0]}\otimes T^{0\rightarrow 1}$ and $A^{-}=A^{-[0]}\otimes T^{0\rightarrow 1}$ and it comes:
$A^{+}=\left(\begin{array}[]{cccccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3&3^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\end{array}\right),\;A^{-}=\left(\begin{array}[]{cccccc}\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&4^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&6^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&1&4^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&2^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\end{array}\right).$
We have $w=(z,\mbox{{\scriptsize$\mathbb{O}$}},y_{2},y_{3},t,h)^{\intercal}$ and the matrices $(A^{\;+},A^{\;-})^{[1]}$ are defined as the result of $\textsf{setrowtozero}(A^{+},A^{-})$:
$A^{+[1]}=\left(\begin{array}[]{cccccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3&3^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right),\;A^{-[1]}=\left(\begin{array}[]{cccccc}\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&4^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&6^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&1&4^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&2^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right).$ (78)
We have $(\mathbbm{y},t)=\{y_{2},y_{3},t\}\neq\emptyset$, $(\mathbbm{y},t)^{+}=\{y_{2}\}$ and $(\mathbbm{y},t)^{\mbox{{\scriptsize$\mathbb{O}$}}}=\{y_{3},t\}$. The mpr is bounded. Indeed, it is easy to check that eg. $I_{\geq}(y_{2})\neq\emptyset$, $I_{\leq}(y_{3})\neq\emptyset$ and $I_{\geq}(t)\neq\emptyset$. The set of dominating variables is $D^{+\mbox{{\scriptsize$\mathbb{O}$}}}=\{y_{2},t\}$. The condition (37) of Theorem 4.2 does not hold and Case $2.2.2$ applies. We have (see 49): $\mathbb{I}^{+\mbox{{\scriptsize$\mathbb{O}$}}}(D)=\{(1,t),(2,2),(4,t)\}$. And the three possibilities are studied in the next array.
$\begin{array}[]{lllllll}\left(\begin{array}[]{c}z\\ t\end{array}\right)_{0,1}&\geq&\left(\begin{array}[]{c}y_{2}\\ 4^{\otimes(-1)}\otimes y_{2}\oplus 6^{\otimes(-1)}\otimes h\end{array}\right)&::&\left(\begin{array}[]{c}\mbox{$[\mathbbm{1}\otimes\lambda,+\infty]$}\\ \mbox{$[6^{\otimes(-1)}\otimes\mu,+\infty]$}\end{array}\right)&::&\left(\begin{array}[]{c}h\mathbb{U}\\ h\mathbb{B}\end{array}\right)\\ \left(\begin{array}[]{c}z\\ y_{2}\end{array}\right)_{0,2}&\geq&\left(\begin{array}[]{c}1\otimes t\oplus 4^{\otimes(-1)}\otimes h\\ 1\otimes t\oplus 4^{\otimes(-1)}\otimes h\end{array}\right)&::&\left(\begin{array}[]{c}\mbox{$[4^{\otimes(-1)}\otimes\mu,+\infty]$}\\ \mbox{$[4^{\otimes(-1)}\otimes\mu,+\infty]$}\end{array}\right)&::&\left(\begin{array}[]{c}h\mathbb{B}\\ h\mathbb{B}\end{array}\right)\\ \left(\begin{array}[]{c}z\\ t\end{array}\right)_{0,4}&\geq&\left(\begin{array}[]{c}y_{2}\\ 3^{\otimes(-1)}\otimes y_{3}\oplus 5^{\otimes(-1)}\otimes h\end{array}\right)&::&\left(\begin{array}[]{c}\mbox{$[\mathbbm{1}\otimes\lambda,+\infty]$}\\ \mbox{$[5^{\otimes(-1)}\otimes\mu,+\infty]$}\end{array}\right)&::&\left(\begin{array}[]{c}h\mathbb{U}\\ h\mathbb{B}\end{array}\right).\end{array}$
By Theorem 4.3: $(\mathbb{T},\mathbb{T}^{\prime})=(\mathbb{B},\mathbb{B})$. So, we have: $(i^{*},j^{*})=(2,2)$ from the previous array which means that $y_{2}=1\otimes t\oplus 4^{\otimes(-1)}\otimes h$.
The set of linear equalities $\mathcal{L}^{[2]}$ is:
$\mathcal{L}^{[2]}=\mathcal{L}^{[1]}\cup\{y_{2}=1\otimes t\oplus 4^{\otimes(-1)}\otimes h\}=\{y_{1}=3^{\otimes(-1)}\otimes h,y_{2}=1\otimes t\oplus 4^{\otimes(-1)}\otimes h\}.$ (79)
The new cost function is:
$\textsf{cost}^{[2]}(y,t,h)=1\otimes t\oplus 4^{\otimes(-1)}\otimes h.$ (80)
We define the $6\times 6$-transition matrix $T^{1\rightarrow 2}$ which corresponds to $y_{2}=1\otimes t\oplus 4^{\otimes(-1)}\otimes h$ as follows:
$T^{1\rightarrow 2}:=\left(\begin{array}[]{cccccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&1&4^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\end{array}\right).$
We calculate the matrices $A^{+}=A^{+[1]}\otimes T^{1\rightarrow 2}$ and $A^{-}=A^{-[1]}\otimes T^{1\rightarrow 2}$ and it comes:
$A^{+}=\left(\begin{array}[]{cccccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&1&4^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3&3^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right),\;A^{-}=\left(\begin{array}[]{cccccc}\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&1&4^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}&6^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&1&4^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&2^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right).$
We have $w=(z,\mbox{{\scriptsize$\mathbb{O}$}},\mbox{{\scriptsize$\mathbb{O}$}},y_{3},t,h)^{\intercal}$ and the matrices $(A^{\;+},A^{\;-})^{[2]}$ are defined as the result of $\textsf{setrowtozero}(A^{+},A^{-})$:
$A^{+[2]}=\left(\begin{array}[]{cccccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3&3^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right),\;A^{-[2]}=\left(\begin{array}[]{cccccc}\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&1&4^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}&6^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&2^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right).\$(81)
$(\mathbbm{y},t)=\{y_{3},t\}\neq\emptyset$ and the mpr problem is bounded because one can check eg. that: $I_{\leq}(y_{3})\neq\emptyset$ and $I_{\geq}(t)\neq\emptyset$. We have $(\mathbbm{y},t)^{+}=\{t\}$ and $(\mathbbm{y},t)^{\mbox{{\scriptsize$\mathbb{O}$}}}=\{y_{3}\}$. The set of dominating variables is $D^{+\mbox{{\scriptsize$\mathbb{O}$}}}=\{t\}$. Indeed, it is easy to see here that $I_{\geq}(y_{3})=\emptyset$ because $A_{[|1,m|]y_{3}}^{+[0]}=\boldsymbol{-\infty}$. Clearly, the condition (37) of Theorem 4.2 does not hold and Case $2.2.2$ applies. We have (see 49): $\mathbb{I}^{+\mbox{{\scriptsize$\mathbb{O}$}}}(D)=I_{\geq}(t)=\{(1,t),(4,t)\}$.
And we have the following two possibilities for the substitution of the dominating variable $t$ which enumerated hereafter:
$\begin{array}[]{lllllll}\left(\begin{array}[]{c}z\\ t\end{array}\right)_{0,1}&\geq&\left(\begin{array}[]{c}4^{\otimes(-1)}\otimes h\\ 6^{\otimes(-1)}\otimes h\end{array}\right)&::&\left(\begin{array}[]{c}\mbox{$[4^{\otimes(-1)}\otimes\mu,+\infty]$}\\ \mbox{$[6^{\otimes(-1)}\otimes\mu,+\infty]$}\end{array}\right)&::&\left(\begin{array}[]{c}h\mathbb{B}\\ h\mathbb{B}\end{array}\right)\\ \left(\begin{array}[]{c}z\\ t\end{array}\right)_{0,4}&\geq&\left(\begin{array}[]{c}3^{\otimes(-1)}\otimes y_{3}\oplus 4^{\otimes(-1)}\otimes h\\ 3^{\otimes(-1)}\otimes y_{3}\oplus 5^{\otimes(-1)}\otimes h\end{array}\right)&::&\left(\begin{array}[]{c}\mbox{$[4^{\otimes(-1)}\otimes\mu,+\infty]$}\\ \mbox{$[5^{\otimes(-1)}\otimes\mu,+\infty]$}\end{array}\right)&::&\left(\begin{array}[]{c}h\mathbb{B}\\ h\mathbb{B}\end{array}\right).\end{array}$
Applying Theorem 4.3: $(\mathbb{T},\mathbb{T}^{\prime})=(\mathbb{B},\mathbb{B})$. We compute the interval $\tau$ and we have: $\tau=[4^{\otimes(-1)}\otimes\mu,+\infty]$. And $\textsf{argMin}(\tau)=\{(1,t),(4,t)\}$ is not a singleton. So, we have to compute $\tau^{\prime}$, Theorem 4.3. And $\tau^{\prime}=[5^{\otimes(-1)}\otimes\mu,+\infty]$. Then, we take $(i^{*},j^{*})\in\textsf{argMin}(\tau^{\prime})=\{(4,t)\}$ which means that $t=3^{\otimes(-1)}\otimes y_{3}\oplus 5^{\otimes(-1)}\otimes h$. We apply the results of § 4.2.3. Then, the set of linear equalities is now:
$\mathcal{L}^{[3]}=\{y_{1}=3^{\otimes(-1)}\otimes h,y_{2}=1\otimes t\oplus 4^{\otimes(-1)}\otimes h\}\cup\{t=3^{\otimes(-1)}\otimes y_{3}\oplus 5^{\otimes(-1)}\otimes h\}.$ (82)
And the new cost function is:
$\textsf{cost}^{[3]}(y,t,h)=3^{\otimes(-1)}\otimes y_{3}\oplus 4^{\otimes(-1)}\otimes h.$ (83)
We define the $6\times 6$-transition matrix $T^{1\rightarrow 2}$ as follows:
$T^{2\rightarrow 3}:=\left(\begin{array}[]{cccccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&5^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\end{array}\right).$
We calculate the matrices $A^{+}=A^{+[2]}\otimes T^{2\rightarrow 3}$ and $A^{-}=A^{-[2]}\otimes T^{2\rightarrow 3}$ and we have:
$A^{+}=\left(\begin{array}[]{cccccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&5^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&2^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right),\;A^{-}=\left(\begin{array}[]{cccccc}\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&4^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&6^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&6^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&5^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&2^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right).$
We have $w=(z,\mbox{{\scriptsize$\mathbb{O}$}},\mbox{{\scriptsize$\mathbb{O}$}},y_{3},\mbox{{\scriptsize$\mathbb{O}$}},h)^{\intercal}$ and the matrices $(A^{\;+},A^{\;-})^{[3]}$ are defined as the result of $\textsf{setrowtozero}(A^{+},A^{-})$:
$A^{+[3]}=\left(\begin{array}[]{cccccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right),\;A^{-[3]}=\left(\begin{array}[]{cccccc}\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&4^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&3^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&5^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right).$ (84)
We have:
$A_{[|1,5|]h}^{+[3]}=\left(\begin{array}[]{c}\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}\\ 3^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right)\geq A_{[|1,5|]h}^{-[3]}=\left(\begin{array}[]{c}\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}\\ 5^{\otimes(-1)}\\ \mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right),$
and the condition (37) of Theorem 4.2 is verified. Then, the function $h\mapsto 4^{\otimes(-1)}$ of the cost function $\textsf{cost}^{[3]}(y,t,h)=3^{\otimes(-1)}\otimes y_{3}\oplus 4^{\otimes(-1)}\otimes h$ is reached at $y_{3}=\mbox{{\scriptsize$\mathbb{O}$}}$. Thus, we have to solve the following triangular linear system of equalities $\mathcal{L}^{[4]}$:
$\begin{array}[]{l}\mathcal{L}^{[4]}:\\ z=3^{\otimes(-1)}\otimes y_{3}\oplus 4^{\otimes(-1)}\otimes h\\ y_{1}=3^{\otimes(-1)}\otimes h\\ y_{2}=1\otimes t\oplus 4^{\otimes(-1)}\otimes h\\ t=3^{\otimes(-1)}\otimes y_{3}\oplus 5^{\otimes(-1)}\otimes h\\ y_{3}=\mbox{{\scriptsize$\mathbb{O}$}}.\end{array}$ (85)
And we have the solution of the mpr obtained by obvious backward substitution in $\mathcal{L}^{[4]}$:
$\min(z)=4^{\otimes(-1)}\otimes h,\;y=\left(\begin{array}[]{c}3^{\otimes(-1)}\otimes h\\ 4^{\otimes(-1)}\otimes h\\ \mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right),\;t=5^{\otimes(-1)}\otimes h.$ (86)
The solution $x$ to the initial fmpr is then: $x=t^{\otimes(-1)}\otimes y=\left(\begin{array}[]{c}2\\ 1\\ \mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right)$ and $x_{2}/3\otimes x_{1}=1\otimes(3\otimes 2)^{\otimes(-1)}=1\otimes 5^{\otimes(-1)}=4^{\otimes(-1)}$. We retrieve the minimum ((4)(-4) with the usual notations for reals) found by the pseudo-polynomial method developed in [10] and applied on Example 3 p. 1469 of [10].
7 Conclusion
In this conclusion we try to provide a more precise analysis of the complexity of the substitution method. We also compare our result with other known results.
For the forward substitution we have to establish the hierarchy between the variables of the problem we need to study $m$ inequalities of the form (3a) which generate at most $n$ AND-inequalities (3c) and at most $n$ OR-inequalities (3b): $\mathcal{O}(2nm)$. From the OR-inequalities the computation complexity of the functions $f^{\geq}_{ij}$ (26a)-(26b)-(26c)-(26d) is $\mathcal{O}(nm)$. From the $f^{\geq}_{ij}$ the computation complexity of the $f_{z,ij}$ functions (the possible next cost function (28)) is also $\mathcal{O}(nm)$. To decide which variable can be substituted we need Theorem 4.3 and compute the intervals $f_{z,ij}(R_{\lambda\mu}(x,h))$ and $f_{ij}(R_{\lambda\mu}(x,h))$ in $\mathcal{O}(2nm)$. This procedure is repeated $n$ times and the overall complexity of the forward susbtitution is thus:
$\mathcal{O}(6n^{2}m).$
Now we have to add the time complexity of the resolution of the triangular system of equalities $\mathcal{L}^{[n]}$ which is $\mathcal{O}(n)$ because we do not need to complute a pseudo inverse of a matrix by residuation theory (see eg. [3]). This is only backward substitution.
Thus, the whole complexity is:
$\mathcal{O}(6n^{2}m)+\mathcal{O}(n).$ (87)
Let us recall that to the best of our knowledge this substitution method is the first strongly polynomial method for such problem. All other proposed schemes of resolution are pseudo-polynomial (see [5], [10], [11]). Thus, because $(\max,+)$-linear programming is now proved to be strongly polynomial then the mean payoff games problem is also strongly polynomial (see eg. [10]). Let us remark that this result was already proved in [19] (english paper) when solving the max-atom problem (MAP). The important consequence of this result is that other six PTIME equivalent problems to MAP which were known to be in NP $\cap$ co-NP (see eg [1], [4], [7], [8], [9], [12], [16], [17], [18], [21]) are also strongly polynonial. We list them hereafter. P1: Looking for non trivial solutions of a tropical cone, P2: Computation of a tropical rank of a matrix, P3: Computation of optimal strategies in parity games (typically: Mean Payoff Games), P4: Scheduling with and/or precedence constraints, P5: Shortest path problem in hypergraph, P6: Model checking and $\mu$-calculus.
And we repeat that the best pseudo-polynomial method known till now can be replaced by our strongly polynomial approach.
Appendix A Maximizing a tropical linear problem
In this appendix we follow the same organization of the main part of the paper dealing with the minimization problem. All the results below are listed without proofs.
To formulate the maximization problem MPR first we borrow the Fourier’s trick (see eg. [13], [20]) used in the usual linear algebra and applying it to the $(\max,+)$-algebra. It comes that $\max(z=c^{\intercal}\otimes x\oplus c_{h}\otimes h)$ is replaced by: $\max(z)\mbox{ and }z\leq c^{\intercal}\otimes x\oplus c_{h}\otimes h$. And then, we will solve the following minimization problem MPR defined as:
$\begin{array}[]{l}\max(z)\end{array}$ (88a)
such that:
$(x,h)\in\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}}^{n+1},$ (88b)
and
$\begin{array}[]{l}z\leq c^{\intercal}\otimes x\oplus c_{h}\otimes h\\ A\otimes x\oplus b\otimes h\leq C\otimes x\oplus d\otimes h.\end{array}$ (88c)
Where the homogenization variable $h$ satisfies the following condition:
$h<+\infty.$ (88d)
And
$\mbox{$z$ and $h$ are not substituable}.$ (88e)
Finally, among the variables of the problem $z,x_{1},\ldots,x_{n}$ we have:
[…]
When necessary such problem will be denoted:
$\textsf{MPR}(A,b,C,d,c,c_{h},z,x,h,(m,n))$.
To the cone $\mathcal{C}(L,W):=\{x:L\otimes x\leq W\otimes x\}$ we associate the function $\overline{\textsf{setrowtozero}}(L,W)$ defined by:
[…]
For the MPR we will consider the following familly of intervals:
$\mathcal{J}:=\{[\mbox{{\scriptsize$\mathbb{O}$}},u],u\in\mathbb{R}_{\mbox{{\scriptsize$\mathbb{O}$}}}\}.$ (90)
Proposition A.1
Let $\mbox{{\scriptsize$\mathbb{O}$}},u_{l}]$, $l=1,\ldots,k$, $k\geq 2$ be $k$ intervals of $\mathcal{J}$ with For all $u_{k}\leq\ldots\leq u_{1}$, we have the following intervals inclusion:
$\mbox{{\scriptsize$\mathbb{O}$}},u_{k}]\subseteq[\mbox{{\scriptsize$\mathbb{O}$}},u_{k-1}]\subseteq\cdots\subseteq[\mbox{{\scriptsize$\mathbb{O}$}},u_{1}].$ (91)
And if $\overline{\textsf{Min}}$ denotes the minimum in the sense of interval inclusion then:
$\cap_{l=1}^{k}[\mbox{{\scriptsize$\mathbb{O}$}},u_{l}]=\overline{\textsf{Min}}\{[\mbox{{\scriptsize$\mathbb{O}$}},u_{l}],l=1,\ldots,k\}.$ (92)
At the step 0 of the substitution the cone associated with the constraints system (6c) of the MPR problem is denoted $\mathcal{C}(A^{\;+},A^{\;-})^{[0]}$ and is defined by:
$\mathcal{C}(A^{+},A^{-})^{[0]}:=\{w\in\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}}^{n+2}:A^{+[0]}\otimes w\leq A^{-[0]}\otimes w\}$ (93a)
where:
$(A^{+},A^{-},w)^{[0]}:=\left(\begin{array}[]{ccc}\mathbbm{1}&\boldsymbol{-\infty}^{\intercal}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \boldsymbol{-\infty}&A&b\end{array}\right),\;\left(\begin{array}[]{ccc}\mbox{{\scriptsize$\mathbb{O}$}}&c^{\intercal}&c_{h}\\ \boldsymbol{-\infty}&C&d\end{array}\right),\;\left(\begin{array}[]{c}z\\ x\\ h\end{array}\right).$ (93b)
We also use the same numerotation convention 4.1 p. 4.1 for matrices and vectors involved in this problem.
The cost function at step 0 of the substitution is denoted and defined by: $\textsf{cost}^{[0]}(x,h):=c^{\intercal}\otimes x\oplus c_{h}\otimes h$, with $c_{h}=\mbox{{\scriptsize$\mathbb{O}$}}$.
The set of the stored linear equalities is denoted $\mathcal{L}^{[k]}$ at each step $k$ of the substitution. And for $k=0$ we have $\mathcal{L}^{[0]}=\emptyset$.
The vector $x$ is called the vector of the remaining variables of the MPR. Of course at step 0 of the method: $x=(x_{j})_{j=1}^{n}$. To $x$ we associate the set of the remaining variables denoted $\mathbbm{x}$.
The parametrized domain of research of a maximum is denoted $\tilde{R}_{\tilde{\lambda}\tilde{\mu}}(x,h)$ and defined by:
$\tilde{R}_{\tilde{\lambda}\tilde{\mu}}(x,h):=\times_{x_{j}\in\mathbbm{x}}[\mbox{{\scriptsize$\mathbb{O}$}},\tilde{\lambda}]\times[\mbox{{\scriptsize$\mathbb{O}$}},\tilde{\mu}$]. (94a)
Where because $h$ must be $\infty$ we can assume that the following condition holds:
$\forall\theta\neq\mbox{{\scriptsize$\mathbb{O}$}},\forall\beta,\;\exists\tilde{\lambda},\tilde{\mu}:\;\beta\otimes\tilde{\mu}<\theta\otimes\tilde{\lambda}$. (94b)
Based on the result of Proposition 3.2 $\forall i=1,\ldots,m$, $\forall j=1,\ldots,n$ such that the following condition is satisfied:
$a^{+}_{i,j}\neq\mbox{{\scriptsize$\mathbb{O}$}}$ and $a^{+}_{i,j}>a^{-}_{i,j}$ (95)
we define the following valid inequality:
$I_{\leq}(x_{j},i):=\{a^{+}_{i,j}\otimes x_{j}\leq a^{-}_{i,.}\otimes w\}.$ (96)
Under condition (95) the inequality $I_{\leq}(x_{j},i)$ can be rewritten as:
$I_{\leq}(x_{j},i)=\{x_{j}\leq f^{\leq}_{ij}(x,h)\}.$ (97)
Where the $(x,h)$-linear function $f^{\leq}_{ij}(x,h)$ is defined by:
$f^{\leq}_{ij}(x,h):=\ell_{ij}(x)\oplus r_{ij}\otimes h,$ (98a)
with $\ell_{ij}$ is the following $x$-linear function which does not depend on $x_{j}$:
$\ell_{ij}(x):=v^{\intercal}_{ij}\otimes x,$ (98b)
where $v_{ij}$ denotes the $n$-dimensional vector such that:
[…]
and the scalar $r_{ij}$ is defined by:
$r_{ij}:=(a^{+}_{i,j})^{\otimes(-1)}\otimes d_{i}.$ (98d)
Assuming that $I_{\leq}(x_{j},i)\neq\emptyset$, or equivalently be a valid inequality, we define the following function:
$f_{z,ij}(x,h):=\oplus_{j^{\prime}\neq j}c_{j^{\prime}}\otimes x_{j^{\prime}}\oplus c_{j}\otimes f^{\leq}_{ij}(x,h)\oplus c_{h}\otimes h.$ (99)
By replacing $f^{\leq}_{ij}(x,h)$ in (27) we obtain the new expression for $f_{z,ij}(x,h)$:
$f_{z,ij}(x,h):=\oplus_{j^{\prime}\neq j}(c_{j^{\prime}}\oplus c_{j}\otimes v_{ij,j^{\prime}})\otimes x_{j^{\prime}}\oplus(c_{h}\oplus c_{j}\otimes r_{ij})\otimes h.$ (100)
The next Theorem will be useful to justify the substitution method developed in this paper.
Theorem A.1 (Saturation of an inequlity)
For all $x_{j}$ such that $I_{\leq}(x_{j},i)$ is valid in the sense of Proposition 3.2 we have:
$\sup_{\{x:\;x_{j}\leq f^{\leq}_{ij}(x,h)\}}\left(\begin{array}[]{c}c^{\intercal}\otimes x\oplus c_{h}\otimes h\\ x_{j}\end{array}\right)\;=\left(\begin{array}[]{c}f_{z,ij}(x,h)\\ f^{\leq}_{ij}(x,h)\end{array}\right).$ (101)
And the supremum is reached at $x$ such that we have:
$x_{j}=f^{\leq}_{ij}(x,h).$ (102)
A.1 Classification of the variables and the linear functions of MPR
Based on the result of Proposition 3.2, $\forall i=1,\ldots,m$, $\forall j=1,\ldots,n$ such that the following condition is satisfied:
$a^{-}_{i,j}\neq\mbox{{\scriptsize$\mathbb{O}$}}$ and $a^{+}_{i,j}<a^{-}_{i,j}$ (103)
we define the following valid inequality:
$I_{\geq}(x_{j},i):=\{a^{-}_{i,j}\otimes x_{j}\geq a^{+}_{i,.}\otimes w\}.$ (104)
Under condition (31) the result of Proposition 3.2 $I_{\geq}(x_{j},i)=\{x_{j}\geq f^{\geq}_{ij}(x,h)\}$ (see (26a)-(26d)).
The ordering between the variables of the problem is denoted $\preceq_{\overline{var}}$ and based on their variation domain such that:
$\{x_{j}\in\mathbbm{x}:\mbox{$I_{\leq}(x_{j})\neq\emptyset$ and $I_{\geq}(x_{j})\neq\emptyset$}\}\preceq_{\overline{var}}\{x_{j}\in\mathbbm{x}:\mbox{$I_{\leq}(x_{j})\neq\emptyset$ and $I_{\geq}(x_{j})=\emptyset$}\}.$ (105)
We also mentioned that:
$\{x_{j}\in\mathbbm{x}:\mbox{$I_{\leq}(x_{j})\neq\emptyset$ and $I_{\geq}(x_{j})=\emptyset$}\}\preceq_{\overline{var}}\{x_{j}\in\mathbbm{x}:\mbox{$I_{\leq}(x_{j})=\emptyset$ and $I_{\geq}(x_{j})=\emptyset$}\}$.
We now define the following partition of non null functions involved in the MPR.
Definition A.1
A non null $(x,h)$-linear function $f:(x,h)\mapsto\alpha^{\intercal}\otimes x\oplus\beta\otimes h$ is said to be $h$-bounded if $\alpha=\boldsymbol{-\infty}$. Otherwise the function $f$ is said to be $h$-unbounded.
[…]
  • $\forall f=\alpha^{\intercal}\otimes x\oplus\beta\otimes h\in h\tilde{\mathbb{U}}$, using interval calculus formulae we have:
    $f(\tilde{R}_{\tilde{\lambda}\tilde{\mu}}(x,h))=[\mbox{{\scriptsize$\mathbb{O}$}},\oplus_{i}\alpha_{i}\otimes\tilde{\lambda}\oplus\beta\otimes\tilde{\mu}$], (106a)
    and by assumption ((94b) with $\theta=\oplus_{i}\alpha_{i}$) on $\tilde{\lambda}$ and $\tilde{\mu}$ we have:
    $\forall\alpha\in\mathbb{R}_{\mbox{{\scriptsize$\mathbb{O}$}}}^{n}\setminus\{\boldsymbol{-\infty}\},\forall\beta:\;f(\tilde{R}_{\tilde{\lambda}\tilde{\mu}}(x,h))=[\mbox{{\scriptsize$\mathbb{O}$}},\oplus_{i}\alpha_{i}\otimes\tilde{\lambda}$] (106b)
    Let us stress that this interval does not depend on $\beta$ and by assumption (94b):
    $\mbox{{\scriptsize$\mathbb{O}$}},\beta\otimes\tilde{\mu}]\subset[\mbox{{\scriptsize$\mathbb{O}$}},\oplus_{i}\alpha_{i}\otimes\tilde{\lambda}$]. (106c)
  • And $\forall f=\beta\otimes h\in h\tilde{\mathbb{B}}$:
    $f(\tilde{R}_{\tilde{\lambda}\tilde{\mu}}(x,h))=[\mbox{{\scriptsize$\mathbb{O}$}},\beta\otimes\tilde{\mu}$]. (107)
From this above results we have the following ordering between the non null linear functions denoted $\preceq_{\overline{fct}}$ which is based on their value domain:
$\{f:f\in h\tilde{\mathbb{B}}\}\preceq_{\overline{fct}}\{f:f\in h\tilde{\mathbb{U}}\}$. (108)
A.2 Choosing a variable for a new substitution after $k$ substitutions in MPR
The procedure for choosing a remaining variable to be substituted is based on the theorems listed below. Recall that we use numerotation convention 4.1 p. 4.1.
Because when $\mathbbm{x}^{+}\neq\emptyset$ the cost function can increases we have the following result.
[…]
The function $h\mapsto c_{h}\otimes h$ is the lower bound of the cost function $(x,h)\mapsto c^{\intercal}_{\mathbbm{x}}\otimes x\oplus c_{h}\otimes h$ which is attained at $x=\boldsymbol{-\infty}$ iff the following conditions hold:
the cost function cannot increase, ie.:
$\mathbbm{x}^{+}=\emptyset,$ (109a)
and the homogenization variable $h$ can take any arbitrary value, ie.:
$A_{[|1,m|]h}^{+[k]}\leq A_{[|1,m|]h}^{-[k]}.$ (109b)
Theorem A.3
Let $\mathbb{I}$ be a subset of valid inequalities $I_{\leq}(x_{j},i)$, ie. a set such that:
$\mathbb{I}\subseteq I_{\leq}(\mathbbm{x}).$ (110)
Where
$I_{\leq}(\mathbbm{x}):=\{(i,j):\mbox{ $x_{j}\in\mathbbm{x}$ and $I_{\leq}(x_{j},i)$ is valid}\}.$ (111)
 $\mbox{ $I_{\leq}(x_{j},i)\neq\emptyset$ or equivalently $I_{\leq}(x_{j},i)$ is valid}\Rightarrow z\leq f_{z,ij}(x,h).$ (112)
Recalling that $f_{z,ij}(x,h)$ is defined by (100).
Let $\mathbb{T}$ and $\mathbb{T}^{\prime}$ denoting either $\tilde{\mathbb{B}}$ or $\tilde{\mathbb{U}}$, respectively. The $2$-tuple $(\mathbb{T},\mathbb{T}^{\prime})$ is a function of $\mathbb{I}$ and is defined according to $\preceq_{fct}$ as follows with priority to the cost function:
[…]
Then, define the following set of indexes:
$\mathcal{T}(\mathbb{T},\mathbb{T}^{\prime},\mathbb{I}):=\{(i,j)\in\mathbb{I}:(f_{z,ij},f^{\leq}_{ij})\in(h\mathbb{T},h\mathbb{T}^{\prime})\}.$ (114)
Let us define the following assertion which deals with the cost $z$ and an interval $\tau$:
$\begin{array}[]{l}\mathrm{Z}(\tau):\\ \forall(i,j)\in\mathcal{T}(\mathbb{T},\mathbb{T}^{\prime},\mathbb{I})\;\forall z\;(z\in\tau\Rightarrow\exists(x,h)\in\tilde{R}_{\tilde{\lambda}\tilde{\mu}}(x,h)\;z\leq f_{z,ij}(x,h)).\end{array}$ (115)
And let us define the following assertion dealing with variables of the problem and intervals $\tau,\tau^{\prime}$:
$\begin{array}[]{l}\mathrm{X}(\tau,\tau^{\prime}):\\ \forall(i,j)\in\mathcal{T}(\mathbb{T},\mathbb{T}^{\prime},\mathbb{I})\cap\overline{\textsf{argMin}}(\tau)\;\forall x_{j}\;(x_{j}\in\tau^{\prime}\Rightarrow\exists(x,h)\in\tilde{R}_{\tilde{\lambda}\tilde{\mu}}(x,h)\;x_{j}\leq f^{\leq}_{ij}(x,h)).\end{array}$ (116)
Then, we have the following equivalences.
$\mathrm{Z}(\tau)\mbox{ true }\Leftrightarrow\tau=\overline{\textsf{Min}}\{f_{z,ij}(\tilde{R}_{\tilde{\lambda}\tilde{\mu}}(x,h)),(i,j)\in\mathcal{T}(\mathbb{T},\mathbb{T}^{\prime},\mathbb{I})\}.$ (117)
And
$\mathrm{X}(\tau,\tau^{\prime})\mbox{ true }\Leftrightarrow\tau^{\prime}=\overline{\textsf{Min}}\{f^{\leq}_{ij}(\tilde{R}_{\tilde{\lambda}\tilde{\mu}}(x,h)),(i,j)\in\overline{\textsf{argMin}}(\tau)\cap\mathcal{T}(\mathbb{T},\mathbb{T}^{\prime},\mathbb{I})\}.$ (118)
We are now in position to indicate how to choose a remaining variable to be substituted. The choice also depends on the following different cases which are listed hereafter.
Case 0: $\max(z)=\mbox{{\scriptsize$\mathbb{O}$}}$ or no maximum. Case $0.1$: $\mathbbm{x}=\emptyset$ and the condition (109b) of Theorem A.2 does not holds. $\mathbbm{x}=\emptyset$ means that $k=n$ (no more variable to be substituted) and $\textsf{cost}^{[n]}(x,h)=c_{h}^{[n]}\otimes h$. The condition (109b) of Theorem A.2 does not holds means $A_{[|1,m|]h}^{+[n]}\nleq A_{[|1,m|]h}^{-[n]}$ then $h=\mbox{{\scriptsize$\mathbb{O}$}}$ is the only possible value for $h$ and by backward substitution we have $z=\mbox{{\scriptsize$\mathbb{O}$}}$, $x=\boldsymbol{-\infty}$.
Case $0.2$: $\mathbbm{x}\neq\emptyset$ and $I_{\leq}(\mathbbm{x})=\emptyset$ and $I_{\geq}(\mathbbm{x})=\emptyset$ and the conditions (109a)-(109b) of Theorem A.2 do not hold. The substitution process stops because no valid inequalities can be generated. No maximum.
Case $1$: switching case, cf. § 2.4. $\mathbbm{x}^{+}=\emptyset$ and $\textsf{cost}^{[k]}(x,h)=c_{h}^{[k]}\otimes h$ and the condition (109b) of Theorem A.2 does not holds and $\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}\neq\emptyset$ and $I_{\leq}(\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}})=\emptyset$ and $I_{\geq}(\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}})\neq\emptyset$. The substitution must switch to the following minimizing problem:
$\textsf{pmr}(A,b,C,d,c,c_{h},z,x,h,(m,n)),$ (119a)
where the matrices $A$ and $C$ are defined by:
$A:=A^{-[k]}_{[|1,m|][|1,n|]},\;C:=A^{+[k]}_{[|1,m|][|1,n|]}.$ (119b)
The vectors $b$ and $d$ are defined by:
$b:=A^{-[k]}_{[|1,m|]h},d:=A^{+[k]}_{[|1,m|]h}$. (119c)
The initial cost function of the MPR is characterized by the vector $c$ and the constant $c_{h}$ such that:
$c:=\boldsymbol{-\infty},\;c_{h}:=c_{h}^{[k]}.$ (119d)
And the variables of the problem are:
$z:=z,\;x\leftarrow x^{\mbox{{\scriptsize$\mathbb{O}$}}},\;h:=h.$ (119e)
Recalling that the notation $x\leftarrow x^{\mbox{{\scriptsize$\mathbb{O}$}}}$ means that we store in the $n$-dimensional vector $x$ the remaining variables of the set $\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}$. The other components (already substituted variables) are set to $\mathbb{O}$.
And the dimensions $(m,n)$ of the mpr are obviously the same as the ones of the MPR.
Case $2$: Not Case 0 and not Case $1$. The first case is Case $2.1$: $\mathbbm{x}=\emptyset$. This means that $k=n$ (no more variable to be substituted) and $\textsf{cost}^{[n]}(x,h)=c_{h}^{[n]}\otimes h$. And if $A_{[|1,m|]h}^{+[n]}\leq A_{[|1,m|]h}^{-[n]}$ then $h$ can take all possible values in particular we can take $h=\mathbbm{1}$ and the MPR succeeds with maximum $z=c_{h}^{[n]}$. Otherwise $h=\mbox{{\scriptsize$\mathbb{O}$}}$ is the only possible value for $h$ and $z=\mbox{{\scriptsize$\mathbb{O}$}}$.
The other case is Case $2.2$: $\mathbbm{x}\neq\emptyset$ and $\exists x_{j}\in\mathbbm{x}$: $I_{\leq}(x_{j})\neq\emptyset$. Let us distinguish the following subcases which are based on the form of the cost function:
$\textsf{cost}^{[k]}(x,h)=c^{\intercal}_{\mathbbm{x}^{+}}\otimes x^{+}\oplus c^{\intercal}_{\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}}\otimes x^{\mbox{{\scriptsize$\mathbb{O}$}}}\oplus c_{h}\otimes h,$ (120a)
where:
$\mbox{$\mathbbm{x}^{+}:=\{x_{j}\in\mathbbm{x}:\;c_{j}>\mbox{{\scriptsize$\mathbb{O}$}}\}$ and $\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}:=\{x_{j}\in\mathbbm{x}:\;c_{j}=\mbox{{\scriptsize$\mathbb{O}$}}\}$}.$ (120b)
And the vectors of remaining $x^{+}$ and $x^{\mbox{{\scriptsize$\mathbb{O}$}}}$ are defined by:
$x^{+}:=(x_{j})_{\{j:x_{j}\in\mathbbm{x}^{+}\}}\mbox{ and }x^{\mbox{{\scriptsize$\mathbb{O}$}}}:=(x_{j})_{\{j:x_{j}\in\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}\}}.$ (120c)
[…]
Case $2.2.2$: otherwise, the conditions (109a)-(109b) of Theorem A.2 do not hold. Based on the hierarchy $\preceq_{\overline{var}}$ (see (105)) between the remaining variables let us define the set of the dominating variables $D^{+\mbox{{\scriptsize$\mathbb{O}$}}}:=\{x_{j}\in\mathbbm{x}:\;I_{\geq}(x_{j})\neq\emptyset\mbox{ and }I_{\leq}(x_{j})\neq\emptyset\}$. And the set of inequalities:
[…]
Where: $I_{\leq}(\mathbbm{x}\cap D^{+\mbox{{\scriptsize$\mathbb{O}$}}}):=\{(i,j):\mbox{ $x_{j}\in\mathbbm{x}\cap D^{+\mbox{{\scriptsize$\mathbb{O}$}}}$ and $I_{\leq}(x_{j},i)$ is valid}\}$.
We apply Theorem A.3 recalling that $(\mathbb{T},\mathbb{T}^{\prime})$ is defined according to $\preceq_{\overline{fct}}$ by (113) and the subset of valid inequalities is $\mathbb{I}:=\mathbb{I}^{+\mbox{{\scriptsize$\mathbb{O}$}}}(D)$. Thus, we are “working” with the indexes of rows and columns (or variables) of the system of inequalities which is: $\mathcal{T}(\mathbb{T},\mathbb{T}^{\prime},\mathbb{I}^{+\mbox{{\scriptsize$\mathbb{O}$}}}(D))$.
Then, we have to compute the interval $\tau$ such that $\mathrm{Z}(\tau)$ is true (see (115) and (117)).
If $\overline{\textsf{argMin}}(\tau)$ is a singleton $\{(i^{*},j^{*})\}$ then take the unique $2$-tuple $(i^{*},j^{*})$ for the substitution of $x_{j^{*}}=f^{\leq}_{i^{*}j^{*}}(x,h)$. Otherwise, take any $2$-tuple $(i^{*},j^{*})$ element of $\overline{\textsf{argMin}}(\tau^{\prime})$ recalling that the interval $\tau^{\prime}$ is the interval such that $\mathrm{X}(\tau,\tau^{\prime})$ is true (see (116) and (118)).
A.3 The new characteristic elements of the MPR after substitution
The new set of equalities is defined as follows. If conditions of Theorem A.2 are verified then set $k+1=n$ and $\mathcal{L}^{[n]}:=\mathcal{L}^{[k]}\cup_{x_{j}\in\mathbbm{x}}\{x_{j}=\mbox{{\scriptsize$\mathbb{O}$}}\}\cup\{z=c_{h}\otimes h\}$. Then, MPR stops and we solve the linear system $\mathcal{L}^{[n]}$ and express all the $x_{j}$’s as function of the homogenization variable $h$. Otherwise, we define $\mathcal{L}^{[k+1]}$ as: $\mathcal{L}^{[k+1]}:=\mathcal{L}^{[k]}\cup\{x_{j^{*}}=f^{\leq}_{i^{*}j^{*}}(x,h)\}$ and we apply Theorem A.1. And based on the previous cases the new cost function is defined as:
$\textsf{cost}^{[k+1]}(x,h):=f_{z,i^{*}j^{*}}(x,h).$ (122)
And in all cases the new cone $\mathcal{C}(A^{+},A^{-})^{[k+1]}$ associated with MPR is deduced from the following set of inequalities:
[…]
and $\ell$ is the $n+2$-dimensional column vector which has the same components as the vector $w^{[k]}$ except its $j^{*}$th component wich is:
$\ell_{j^{*}}:=f^{\leq}_{i^{*}j^{*}}(x,h).$ (123b)
Let us define the following $(n+2)\times(n+2)$ transition matrix $T^{k\rightarrow k+1}$ by:
$\forall j:t^{k\rightarrow k+1}_{j,.}:=\left\{\begin{array}[]{lc}e^{\intercal}_{j}\mbox{ if $j\neq j^{*}$}\\ (\mbox{{\scriptsize$\mathbb{O}$}},v^{\intercal}_{i^{*}j^{*}},r_{i^{*}j^{*}})&\mbox{ if $j=j^{*}$.}\end{array}\right$. (124)
Where $e^{\intercal}_{j}$ denotes the $j$th $n+2$-dimensional row vector of the $(n+2)\times(n+2)$-identity matrix $\boldsymbol{I}_{n+2}$.
The new matrices $(A^{+},A^{-})^{[k+1]}$ are defined by:
$(A^{+},A^{-})^{[k+1]}:=\overline{\textsf{setrowtozero}}(A^{+[k]}\otimes T^{k\rightarrow k+1},A^{-[k]}\otimes T^{k\rightarrow k+1})$ (125)
recalling that function $\overline{\textsf{setrowtozero}}$ is defined by (89).
And the new vector $w^{[k+1]}$ is defined by:
[…]
Finally, the new set of the remaining variables at step $k+1$ is defined by:
$\mathbbm{x}:=\mathbbm{x}\setminus\{x_{j^{*}}\}.$ (127)
A.4 A numerical example
In this section we illustrate our strongly polynomial method on a numerical example borrowed from the litterature which are solved using pseudo-polynomial method in [10]. And we apply substitution on [10], Example $1$ p. 1458 which is:
$\max\{z=1\otimes x_{1}\oplus 3\otimes x_{2},\;x=(x_{1},x_{2})^{\intercal}\in\mathcal{P}(A,b,C,d)\cup\{\boldsymbol{\pm\infty}\}\},$
where:
$A:=\left(\begin{array}[]{cc}\mbox{{\scriptsize$\mathbb{O}$}}&1^{\otimes(-1)}\\ 2^{\otimes(-1)}&2^{\otimes(-1)}\\ 1^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right),\;b:=\boldsymbol{-\infty},\;C:=\left(\begin{array}[]{cc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&2\end{array}\right),\;d:=\left(\begin{array}[]{c}\mathbbm{1}\\ \mathbbm{1}\\ \mathbbm{1}\\ \mathbbm{1}\end{array}\right).$
With the following change of notations $c\leftrightarrow b$ and $B\leftrightarrow C$.
The MPR is defined on the homogenized cone $\mathcal{C}(A,b,C,d)$ by adding the homogenization variable $h$.
At step 0 of the substitution the cone associated with the above constraints is denoted $\mathcal{C}(A^{+},A^{-})^{[0]}$ and is characterized by the following system of inequalities:
$A^{+[0]}\otimes w\leq A^{-[0]}\otimes w,$ (128a)
with $w=(z,x_{1},x_{2},h)^{\intercal}\in\overline{\mathbb{R}}^{4}_{\mbox{{\scriptsize$\mathbb{O}$}}}$ and the matrices $(A^{+},A^{-})^{[0]}$ are defined by:
$A^{+[0]}:=\left(\begin{array}[]{cccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&1^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&2^{\otimes(-1)}&2^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&1^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right),\;A^{-[0]}:=\left(\begin{array}[]{cccc}\mbox{{\scriptsize$\mathbb{O}$}}&1&3&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&2&\mathbbm{1}\end{array}\right).$ (128b)
And the Fourier’s trick for this example is:
$\max(z)$ (129a)
$z\leq\textsf{cost}^{[0]}(x,h)$ (129b)
$\textsf{cost}^{[0]}(x,h)=1\otimes x_{1}\oplus 3\otimes x_{2}\oplus\mbox{{\scriptsize$\mathbb{O}$}}\otimes h.$ (129c)
$\mathbbm{x}=\{x_{1},x_{2}\}$. It is easy to check that $\forall x_{j}\in\mathbbm{x}:I_{\leq}(x_{j})\cup I_{\geq}(x_{j})\neq\emptyset$. So that the MPR is bounded. We have $\mathbbm{x}^{+}=\{x_{1},x_{2}\}$ and $\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}=\emptyset$. The set of dominating variables is $D^{+\mbox{{\scriptsize$\mathbb{O}$}}}=\{x_{1},x_{2}\}$. The condition (37) of Theorem 4.2 is not verified. Thus, Case $2.2.2$ applies. We have (see (121)): $\mathbb{I}^{+\mbox{{\scriptsize$\mathbb{O}$}}}(D)=I_{\leq}(\{x_{1},x_{2}\})=\{(1,2),(2,1),(2,2),(3,1),(4,1)\}$.
And we enumerate all the possibilities in the following array:
[…]
The set $\tau$, Theorem A.3, is such that $\tau=[\mbox{{\scriptsize$\mathbb{O}$}},1\otimes\tilde{\lambda}$]. And $\overline{\textsf{argMin}}(\tau)=\{(2,2)\}$ which is a singleton. Thus, $(i^{*},j^{*})=(2,2)$. So, we have: $x_{2}=2\otimes h$ in row $2$ of the linear system of inequalities (128a)-(128b).
Applying results of § A.3 the new characteristic elements are listed below.
$\mathcal{L}^{[1]}=\{x_{2}=2\otimes h\}.$ (130)
The new cost function is:
$\textsf{cost}^{[1]}(x,h):=f_{z,22}(x,h)=1\otimes x_{1}\oplus 5\otimes h.$ (131)
The $4\times 4$-transition matrix is defined by:
$T^{0\rightarrow 1}:=\left(\begin{array}[]{cccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&2\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\end{array}\right).$ (132)
Then, we compute the following matrices $A^{+}:=A^{+[0]}\otimes T^{0\rightarrow 1}$ and $A^{-}:=A^{-[0]}\otimes T^{0\rightarrow 1}$ and we obtain:
$A^{+}:=\left(\begin{array}[]{cccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&1\\ \mbox{{\scriptsize$\mathbb{O}$}}&2^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&1^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\end{array}\right),\;A^{-}:=\left(\begin{array}[]{cccc}\mbox{{\scriptsize$\mathbb{O}$}}&1&\mbox{{\scriptsize$\mathbb{O}$}}&5\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&2\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&4\end{array}\right).$
The new cone $\mathcal{C}(A^{+},A^{-})^{[1]}$ is defined by:
$A^{+[1]}\otimes w\leq A^{-[1]}\otimes w,$ (133a)
with $w=(z,x_{1},\mbox{{\scriptsize$\mathbb{O}$}},h)^{\intercal}\in\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}}^{4}$ and the matrices $(A^{\;+},A^{\;-})^{[1]}$ are defined as a result of $\overline{\textsf{setrowtozero}}(A^{+},A^{-})$:
$A^{+[1]}:=\left(\begin{array}[]{cccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&1\\ \mbox{{\scriptsize$\mathbb{O}$}}&2^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&1^{\otimes(-1)}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\end{array}\right),\;A^{-[1]}:=\left(\begin{array}[]{cccc}\mbox{{\scriptsize$\mathbb{O}$}}&1&\mbox{{\scriptsize$\mathbb{O}$}}&5\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&2\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&4\end{array}\right).$ (133b)
$\mathbbm{x}=\{x_{1}\}\neq\emptyset$. It is easy to see that the MPR problem is bounded. We have $\mathbbm{x}^{+}=\{x_{1}\}$ and $\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}=\emptyset$. The set of dominating variables is $D^{+\mbox{{\scriptsize$\mathbb{O}$}}}:=\{x_{1}\}$. The conditions (109a)-(109b) of Theorem A.2 are not verified. Thus, Case $2.2.2$ applies. We have (see (121)): $\mathbb{I}^{+\mbox{{\scriptsize$\mathbb{O}$}}}(D)=I_{\leq}(\{x_{1}\})=\{(2,1),(3,1),(4,1)\}$.
And we enumerate all the possibilities in the following array:
[…]
Applying Theorem A.3 we have: ($𝔹~)(\mathbb{T},\mathbb{T}^{\prime}$)=($\tilde$$\mathbb$$\tilde{\mathbb{B}}$). And the set $\tau$ is $\tau=[\mbox{{\scriptsize$\mathbb{O}$}},5\otimes\tilde{\mu}$]. We remark that $\overline{\textsf{argMin}}(\tau)=\{(2,1),(3,1),(4,1)\}$ is not a singleton. So, we have to compute $\tau^{\prime}$, Theorem A.3, $\tau^{\prime}=[\mbox{{\scriptsize$\mathbb{O}$}},2\otimes\tilde{\mu}$]. And we take $(i^{*},j^{*})\in\overline{\textsf{argMin}}(\tau^{\prime})=\{(2,1)\}$.
Thus, $(i^{*},j^{*})=(2,1)$. So, we have: $x_{1}=2\otimes h$ in row $2$ of the linear system of inequalities (133a)-(133b).
Applying results of § A.3 the new characteristic elments of the problem are listed below.
$\mathcal{L}^{[2]}=\mathcal{L}^{[1]}\cup\{x_{1}=2\otimes h\}=\{x_{2}=2\otimes h,x_{1}=2\otimes h\}.$ (134)
The new cost function is:
$\textsf{cost}^{[2]}(x,h):=f_{z,21}(x,h)=5\otimes h.$ (135)
The $4\times 4$-transition matrix is defined by:
$T^{1\rightarrow 2}:=\left(\begin{array}[]{cccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&2\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\end{array}\right).$ (136)
We compute the following matrices $A^{+}:=A^{+[1]}\otimes T^{1\rightarrow 2}$ and $A^{-}:=A^{-[1]}\otimes T^{1\rightarrow 2}$ and we obtain:
$A^{+}:=\left(\begin{array}[]{cccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&1\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&1\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&2\end{array}\right),\;A^{-}:=\left(\begin{array}[]{cccc}\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&5\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&2\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mathbbm{1}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&2\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&4\end{array}\right).$
The new cone $\mathcal{C}(A^{+},A^{-})^{[2]}$ is defined by:
$A^{+[2]}\otimes w\leq A^{-[2]}\otimes w,$ (137a)
with $w=(z,x_{1},\mbox{{\scriptsize$\mathbb{O}$}},h)^{\intercal}\in\overline{\mathbb{R}}_{\mbox{{\scriptsize$\mathbb{O}$}}}^{4}$ and the matrices $(A^{\;+},A^{\;-})^{[2]}$ are defined as a result of $\overline{\textsf{setrowtozero}}(A^{+},A^{-})$:
$A^{+[2]}:=\left(\begin{array}[]{cccc}\mathbbm{1}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right),\;A^{-[2]}:=\left(\begin{array}[]{cccc}\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&5\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\\ \mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}&\mbox{{\scriptsize$\mathbb{O}$}}\end{array}\right).$ (137b)
Clearly, $A_{[|1,4|]h}^{+[2]}=\boldsymbol{-\infty}\leq A_{[|1,4|]h}^{-[2]}=\boldsymbol{-\infty}$. And the conditions (109a)-(109b) of Theorem A.2 are verified and the cost function $h\mapsto 5\otimes h$ is reached.
The solution of the MPR is obtained by a trivial backward substitution:
$z=5\otimes h,\;x_{2}=2\otimes h,\;x_{1}=2\otimes h.$
And we retrieve by our strongly polynomial metyhod the same maximum as in Example 1 p. 1458 of [10]. But the corner $(2,2)$ is different from the one obtained by the pseudo-polynomial method developed by the authors of [10], and mentioned p. 1474: $x_{1}=1,x_{2}=2$.